### v1

# On determinantal point processes with nonsymmetric kernels

On determinantal point processes with nonsymmetric kernels
Arnaud Poinas
Laboratoire de Mathematiques et Applications, University of Poitiers, Poitiers, France
arnaud.poinas@univ-poitiers.fr
(June 5, 2024)
In $1975$, Odile Macchi [20] introduced determinantal point processes (or DPPs for short) to model fermion particles. Since then, DPPs have seen a wide range of applications not only in physics but also in random matrix theory [24] to model the distributions of eigenvalues of some classes of random matrices, in statistics [18] to model datasets of repulsive point patterns and in machine learning [16] to select a diverse subset of a large dataset, just to name a few examples. In almost all cases, it is assumed that the determinantal point process possess a symmetric kernel which is an important assumption for a lot of results (well-definedness, simulation, distribution of the number of points, …). Although, some examples of DPPs with nonsymmetric kernels are known [3] and there has been a few papers working with them [8, 10], usually for machine learning applications. Nevertheless, previous works only focused on some specific questions related to these DPPs but does not establish their general properties. Our goal in this paper is therefore to investigate how the usual results on DPPs are modified when considering a nonsymmetric kernel. In particular, we are interested in using nonsymmetric kernels in order to construct attractive coupling of DPPs to model marked point processes with repulsion between points of the same mark and attraction between points of different marks. One such example is the ants dataset [11] in the spatstat R package [2] containing locations of nests of two species of ants. While ants of the same species tends to spread out their nests to avoid having to share resources, hence the repulsion in their nest location, it is suspected that one of the two species purposefully put their nest close to one of the other species since they eat dead insects, hence the attraction between the location of nests of different species. While we can use DPPs with symmetric kernels to model the location of the nest of each species, the natural way to create couplings of such DPPs (see Section 1.2) only allows for negative cross correlations hence the interest in using non symmetric kernels in order to allow positive cross correlations between the two DPPs.
In Section 1 we recall the basic definitions of DPPs over a finite set. In Section 2 we use the known results on $P_{0}$ matrices to get necessary and sufficient conditions for a DPP to be well-defined. In Section 3 we show some particular examples of DPP couplings. In Section 4 we present an attempt to generalize the mixing property of DPPs which is used in various cases, the most common one being for their simulation. Finally, in Section 5 we show an example of construction and simulation of an attractive coupling of repulsive DPPs.
1 Definitions and first properties of DPPs
1.1 Notations and definitions
[…]
$M\preccurlyeq N\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ N-M\in\mathcal{S}^{+}_{n}(\mathbb{R}).$
We write $I_{n}$ for the $n\times n$ identity matrix and $1_{n}$ (resp. $0_{n}$) for the vector of $\mathbb{R}^{n}$ uniquely composed of $1$s (resp. 00s). For any vector $x=(x_{1},\cdots,x_{n})^{T}\in\mathbb{R}^{n}$ we write $D(x)$ for the $n\times n$ diagonal matrix whose diagonal elements are $x_{1},\cdots,x_{n}$. Thus, $I_{n}=D(1_{n})$.
We first recall the general definition of determinantal measures and $L$-ensemble measures.
Definition 1.1.
Let $\0n\in\mathbb{N}\backslash\{0\}$.
  • Let $L\in\mathcal{M}_{n}(\mathbb{R})$ such that $I_{n}+L$ is invertible. The L-ensemble measure $\mu$ on $\mathcal{P}([n])$ with kernel $L$ is defined by
    $\forall S\subset[n],\leavevmode\nobreak\ \mu(S)=\frac{\det(L_{S})}{\det(I_{n}+L)}.$
  • Let $K\in\mathcal{M}_{n}(\mathbb{R})$. The determinantal measure $\mu$ on $\mathcal{P}([n])$ with kernel $K$ is defined by
    $\forall S\subset[n],\leavevmode\nobreak\ \mu(\{X\in\mathcal{P}([n])\leavevmode\nobreak\ \mbox{s.t.}\leavevmode\nobreak\ S\subset X\})=\det(K_{S}).$
As shown in [16] these two definitions are almost equivalent since, when $I_{n}-K$ is invertible, a determinantal measure with kernel $K$ is an $L$-ensemble measure with kernel $L=K(I_{n}-K)^{-1}$. Reciprocally, an $L$-ensemble measure with kernel $L$ is a determinantal measure with kernel $K=L(I_{n}+L)^{-1}$. These two measures are always well defined and with total mass equal to $1$ as a consequence of the identity
$\forall M\in\mathcal{M}_{n}(\mathbb{R}),\leavevmode\nobreak\ \sum_{S\subset[n]}\det(M_{S})=\det(I_{n}+M).$
Therefore, these measures are probability measures if and only if $\mu(S)\geqslant 0$ for all $S\subset[n]$. The associated probability distributions are called $L$-ensemble and determinantal point processes.
Definition 1.2.
Let $\0n\in\mathbb{N}\backslash\{0\}$.
  • An L-ensemble with kernel $L\in\mathcal{M}_{n}(\mathbb{R})$ is defined as a random variable $X\in\mathcal{P}([n])$ such that
    $\forall S\subset[n],\leavevmode\nobreak\ \mathbb{P}(X=S)=\frac{\det(L_{S})}{\det(I_{n}+L)}.$
    If this probability distribution is well-defined we write $X\sim DPP_{L}(L)$.
  • A determinantal point process (DPP) with kernel $K\in\mathcal{M}_{n}(\mathbb{R})$ is defined as a random variable $X\in\mathcal{P}([n])$ such that
    $\forall S\subset[n],\leavevmode\nobreak\ \mathbb{P}(S\subset X)=\det(K_{S}).$
    If this probability distribution is well-defined we say that $K$ is a DPP kernel and we write $X\sim DPP(K)$.
There are no general results on which conditions are needed for a matrix $K$ to be a DPP kernel. Given a matrix $L$, it is easy to see that its associated L-ensemble measure is a probability distribution if and only if its principal minors are non-negative. A matrix satisfying this property is called a $P_{0}$ matrix.
Proposition 1.3.
The $L$-ensemble measure with kernel $L\in\mathcal{M}_{n}(\mathbb{R})$ is a probability measure if and only if $L$ is a $P_{0}$ matrix meaning that
$\forall S\subset[n],\leavevmode\nobreak\ \leavevmode\nobreak\ \det(L_{S})\geqslant 0.$
Proof.
We already mentioned that the $L$-ensemble measure is a probability measure if and only if
$\forall S\subset[n],\leavevmode\nobreak\ \leavevmode\nobreak\ \frac{\det(L_{S})}{\det(I_{n}+L)}\geqslant 0.$ (1.1)
Taking $S=\emptyset$ shows that we need $\det(I_{n}+L)$ to be positive hence (1.1) is equivalent to $\det(L_{S})\geqslant 0$ for all $S\subset[n]$ and $\det(I_{n}+L)\geqslant 0$. But, since $\det(I_{n}+L)=\sum_{S\subset[n]}\det(L_{S})$ then (1.1) is simply equivalent to $\det(L_{S})\geqslant 0$ for all $S\subset[n]$. ∎
$P_{0}$ matrices were first studied in [7] and we refer to [14] for an extensive review of their properties. We finish by mentioning that when $K$ and $L$ are symmetric matrices, which is the usual setting of determinantal point processes, then it is well known which assumptions are needed on $K$ and $L$ to get probability measures (see [16] for example).
Proposition 1.4.
Let $K,L\in\mathcal{S}_{n}(\mathbb{R})$. $K$ is a DPP kernel if and only if its eigenvalues are in $[0,1]$. The $L$-ensemble measure with kernel $L$ is a probability measure if and only if $L\in\mathcal{S}_{n}^{+}(\mathbb{R})$.
1.2 DPP couplings
There is a natural way of making couplings of determinantal point processes. To express such couplings we begin by pointing out that a coupling of two point processes on $[n]$ can be seen as a point process on $[2n]$ using the bijection
$\begin{array}[]{ccc}[2n]&\leftrightarrow&[n]\times[n]\\ X&\leftrightarrow&(X_{1},X_{2})\end{array}\leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \left\{\begin{array}[]{l}X_{1}=X\cap[n];\\ X_{2}=\big{\{}i-n,i\in X\cap\{n+1,\cdots,2n\}\big{\}}.\end{array}\right.$
As an abuse of notation, we will often write a DPP on $[2n]$ as $(X_{1},X_{2})$ using this bijection. Now, we consider two DPP kernels $K^{(1)},K^{(2)}\in\mathcal{M}_{n}(\mathbb{R})$ and a $2\times 2$ block matrix of the form
$\mathbb{K}=\begin{pmatrix}K^{(1)}&M\\ N&K^{(2)}\end{pmatrix}\in\mathcal{M}_{2n}(\mathbb{R})$
such that $\mathbb{K}$ is a DPP kernel. If $(X_{1},X_{2})\sim DPP(\mathbb{K})$ then $X_{1}\sim DPP(K_{1})$ and $X_{2}\sim DPP(K_{2})$. This way, we can write a coupling of two DPPs on $[n]$ as a DPP on $[2n]$. Such a setting has been considered before [9] but only with symmetric kernels. When $\mathbb{K}$ is symmetric then
$\mathbb{P}(i\in X_{1},j\in X_{2})-\mathbb{P}(i\in X_{1})\mathbb{P}(j\in X_{2})=-\mathbb{K}_{i,j+n}^{2}\leqslant 0$
and therefore this coupling only allows negative cross-correlations. This is why we suggest the possibility of using non-symmetric matrices in order to keep this very natural construction of DPP coupling while allowing some positive cross-dependencies.
1.3 The cardinal of DPPs
Just like in the symmetrical case, the moments of the number of points of a DPP only depends on the set of eigenvalues (taken with multiplicities) of $K$ (and thus $L$). We recall the proof below which does not use the assumption that $K$ is symmetric.
Proposition 1.5.
Let $X\sim DPP(K)$ with $K\in\mathcal{M}_{n}(\mathbb{R})$. We denote the complex eigenvalues of $K$ (taken with multiplicities) by $\lambda_{1},\cdots,\lambda_{n}$. Then,
$\mathbb{E}\left[\binom{|X|}{k}\right]=e_{k}(\lambda_{1},\cdots,\lambda_{n}),$
where $e_{k}$ is the $k$-th elementary symmetric polynomial.
Proof.
Since $k!\binom{|X|}{k}$ is the number of $k$-tuples of elements of $|X|$ then
$k!\binom{|X|}{k}=\sum_{i_{1},\cdots,i_{k}\in[n]}^{\neq}\mathbbm{1}_{i_{1},\cdots,i_{k}\in X}$
hence
$\mathbb{E}\left[k!\binom{|X|}{k}\right]=\sum_{i_{1},\cdots,i_{k}\in[n]}^{\neq}\mathbb{P}(i_{1},\cdots,i_{k}\in X)=k!\sum_{\begin{subarray}{c}S\subset[n]\\ |S|=k\end{subarray}}\mathbb{P}(S\subset X)=k!\sum_{\begin{subarray}{c}S\subset[n]\\ |S|=k\end{subarray}}\det(K_{S}).$
It is well-known (see [12, Theorem 1.2.16] for example) that the sum of minors of size $k$ of $K$ is equal to $e_{k}(\lambda_{1},\cdots,\lambda_{n})$ hence the result. ∎
As a first consequence, we get a simple expression for the expectation and the variance of the number of points of a DPP.
Corollary 1.6.
Let $X\sim DPP(K)$. Then
$\mathbb{E}[|X|]=\textnormal{Tr}(K)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \textnormal{Var}(|X|)=\textnormal{Tr}(K)-\textnormal{Tr}(K^{2}).$
As a second consequence, we get that two DPP kernels with the same eigenvalues may not yields the same distribution on $\mathcal{P}([n])$ but their number of points will have the same distribution.
Corollary 1.7.
Let $X\sim DPP(K)$ and $Y\sim DPP(K^{\prime})$ such that $K$ and $K^{\prime}$ have the same eigenvalues taken with multiplicities. Then, $|X|$ and $|Y|$ have the same distribution.
1.4 The particle-hole involution
One of the earliest known way to create DPPs with nonsymmetric kernels is to use what is often called the particle-hole transformation. For any subset $S$ of $[n]$, the particle-hole transformation is defined by the involution $X\mapsto(X\cap S^{c})\cup(X^{c}\cap S)$ that switches the states of the points in and out of $X\cap S$. It is shown in [4] that determinantal distributions are stable by this operation.
Proposition 1.8 ([4]).
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel and let $X\sim DPP(K)$. For a given set $S\subset[n]$ we define
$\widetilde{X}=(X\cap S^{c})\cup(X^{c}\cap S).$
Then, $\widetilde{X}\sim DPP(\widetilde{K})$ with
$\widetilde{K}=D(\mathbbm{1}_{S})(I_{n}-K)+D(\mathbbm{1}_{S^{c}})K.$
After some permutations of the rows and columns of $K$ and $\widetilde{K}$ we can write
$K=\begin{pmatrix}K_{S}&K_{S,S^{c}}\\ K_{S^{c},S}&K_{S^{c}}\end{pmatrix}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \widetilde{K}=\begin{pmatrix}I_{|S|}-K_{S}&-K_{S,S^{c}}\\ K_{S^{c},S}&K_{S^{c}}\end{pmatrix}.$
As a direct consequence we get an expression of all probabilities $\mathbb{P}(X=S)$ using only the kernel $K$:
Corollary 1.9 (Identity (147) in [16]).
$\mathbb{P}(X=S^{c})=\mathbb{P}([n]\subset\widetilde{X})=\det(D(\mathbbm{1}_{S})(I_{n}-K)+D(\mathbbm{1}_{S^{c}})K)=\det\begin{pmatrix}I_{|S^{c}|}-K_{S^{c}}&-K_{S^{c},S}\\ K_{S,S^{c}}&K_{S}\end{pmatrix}.$
We now give a generalization of Proposition 1.8 where the state of each point in $[n]$ is switched independently with its own probability.
Proposition 1.10.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel and let $X\sim DPP(K)$. Let $p_{1},\cdots,p_{n}\in[0,1]$ and $B_{i}\sim b(p_{i})$ be independent random variables. We define $\widetilde{X}$ as
$\widetilde{X}=\{i\in[n]\leavevmode\nobreak\ \mbox{s.t.}\leavevmode\nobreak\ i\in X\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B_{i}=0\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ i\notin X\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B_{i}=1\}.$
In other word, when $B_{i}=1$ we switch the state of $i$ within $X$. Then,
$X\sim DPP(D(p)(I_{n}-K)+D(1_{n}-p)K)$
Proof.
By conditioning with respect to the $B_{i}$ we get
$\mathbb{P}(S\subset\widetilde{X})$ $=\sum_{T\subset S}\mathbb{P}(S\subset\widetilde{X}|B_{i}=1\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ i\in T\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B_{i}=0\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ i\in T\backslash S)\prod_{i\in T}p_{i}\prod_{i\in S\backslash T}(1-p_{i})$
$=\sum_{T\subset S}\mathbb{P}(S\backslash T\subset X,X\cap T=\emptyset)\prod_{i\in T}p_{i}\prod_{i\in S\backslash T}(1-p_{i}).$
By using the particle-hole transform on $X\cap S$ for points in $T$ we get
$\mathbb{P}(S\subset\widetilde{X})$ $=\det\begin{pmatrix}I_{|T|}-K_{T}&-K_{T,S\backslash T}\\ K_{S\backslash T,T}&K_{S\backslash T}\end{pmatrix}\prod_{i\in T}p_{i}\prod_{i\in S\backslash T}(1-p_{i})$
$=\sum_{T\subset S}\det\left(\begin{pmatrix}D(p)_{T}&0\\ 0&D(1_{n}-p)_{S\backslash T}\end{pmatrix}\begin{pmatrix}I_{|T|}-K_{T}&-K_{T,S\backslash T}\\ K_{S\backslash T,T}&K_{S\backslash T}\end{pmatrix}\right)$
Since $D(p)$ and $D(1_{n}-p)$ are diagonal then the matrix inside the determinant in the last identity has its rows indexed by $T$ corresponding to the rows of $(D(p)(I_{n}-K))_{S}$ and its rows indexed by $S\backslash T$ corresponding to the rows of $(D(1_{n}-p)(K))_{S}$. Therefore, by the multilinearity of the determinant with respect to its rows we get that
$\mathbb{P}(S\subset\widetilde{X})=\det\big{(}(D(p)(I_{n}-K)+D(1_{n}-p)K)_{S}\big{)},$
concluding the proof. ∎
We retrieve the usual particle-hole involution result (Proposition 1.8) by taking $p_{i}=\mathbbm{1}_{i\in S}$ for some $S\subset[n]$. We now show that the particle hole involution is closely linked to a common tool appearing in the theory of $P_{0}$ matrices called the principal pivot transform.
Definition 1.11.
Let $M\in\mathcal{M}_{n}(\mathbb{C})$ and $S\subset[n]$ such that $M_{S}$ is invertible. With the right permutation of rows and columns we can write $M$ as $\begin{pmatrix}M_{S}&M_{S,S^{c}}\\ M_{S^{c},S}&M_{S^{c}}\end{pmatrix}.$ The principal pivot transform of $M$ relative to $S$ is then defined as
$ppt(M,S):=\begin{pmatrix}M_{S}^{-1}&-M_{S}^{-1}M_{S,S^{c}}\\ M_{S^{c},S}M_{S}^{-1}&M_{S^{c}}-M_{S^{c},S}M_{S}^{-1}M_{S,S^{c}}\end{pmatrix}.$
The definition of the principal pivot transform is often better understood through the following result.
Proposition 1.12.
Let $M\in\mathcal{M}_{n}(\mathbb{C})$ and $S\subset[n]$ such that $M_{S}$ is invertible and $x,y\in\mathbb{R}^{n}$. We use the same permutation as in Definition 1.11 ans write $x=\begin{pmatrix}x_{S}\\ x_{S^{c}}\end{pmatrix}$ and $y=\begin{pmatrix}y_{S}\\ y_{S^{c}}\end{pmatrix}$. Then,
$\begin{pmatrix}y_{S}\\ y_{S^{c}}\end{pmatrix}=M\begin{pmatrix}x_{S}\\ x_{S^{c}}\end{pmatrix}\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \begin{pmatrix}x_{S}\\ y_{S^{c}}\end{pmatrix}=ppt(M,S)\begin{pmatrix}y_{S}\\ x_{S^{c}}\end{pmatrix}.$
Now, we can link the principal pivot transform and DPPs by proving the following result.
Proposition 1.13.
Let $L\in\mathcal{M}_{n}(\mathbb{R})$ be a $P_{0}$ matrix such that $L_{S}$ is invertible and let $X\sim DPP_{L}(L)$. For a given set $S\subset[n]$ we define the particle-hole transformation of $X$ as
$\widetilde{X}=(X\cap S^{c})\cup(X^{c}\cap S).$
If $I_{n}-\widetilde{K}$ is invertible then $\widetilde{X}\sim DPP_{L}(ppt(L,S))$.
Proof.
Let $K$ (resp. $\widetilde{K}$) the DPP kernel of $X$ (resp. $\widetilde{X}$). After some row and column permutations and using Proposition 1.8 we can write
$K=\begin{pmatrix}K_{S}&K_{S,S^{c}}\\ K_{S^{c},S}&K_{S^{c}}\end{pmatrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \widetilde{K}=\begin{pmatrix}I_{|S|}-K_{S}&-K_{S,S^{c}}\\ K_{S^{c},S}&K_{S^{c}}\end{pmatrix}.$
Since $I_{n}-\widetilde{K}$ is invertible we can define $\widetilde{L}=(I_{n}-\widetilde{K})^{-1}-I_{n}$. We then choose $x,y\in\mathbb{R}^{n}$ such that $y=\widetilde{L}x$ indexed as in Proposition 1.13. Then,
$y=\widetilde{L}x\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ (x+y)=(I_{n}+\widetilde{L})x\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ (I_{n}-\widetilde{K})(x+y)=x\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ y=\widetilde{K}(x+y).$
This expression can be extended into
$\begin{pmatrix}y_{S}\\ y_{S^{c}}\end{pmatrix}=\begin{pmatrix}x_{S}+y_{S}-K_{S}(x_{S}+y_{S})-K_{S,S^{c}}(x_{S^{c}}+y_{S^{c}})\\ K_{S^{c},S}(x_{S}+y_{S})+K_{S^{c}}(x_{S^{c}}+y_{S^{c}})\end{pmatrix}$
[…]
By Proposition 1.12 we can conclude that $\widetilde{L}=ppt(L,S)$.
The result that determinantal measures are stable by particle-hole transformations is thus almost equivalent to the result that $P_{0}$ matrices are stable by principal pivot transforms [14, Theorem 4.8.4]. This illustrates the close link there is between the theory of $P_{0}$ matrices and the theory of DPPs with nonsymmetric kernels.
2 Properties of DPPs with nonsymmetric kernel
2.1 Characterization of DPP kernels
Checking whether a matrix $L$ is an $L$-ensemble kernel or not is equivalent to checking whether or not it is a $P_{0}$ matrix. There is already a lot of literature on the subject and we refer to [14, Section 4.6] for an overview. We focus instead on adapting these results on $P_{0}$ matrices to characterize whether a given matrix $K$ is a DPP kernel or not. We first recall that it was proven in [6] that the problem of checking if a given matrix is a $P$ matrix (matrix with positive principal minors) is co-NP-complete. The same can be expected for checking if a given matrix is a DPP kernel since we can go from $K$ to $L$ in polynomial time. We thus can’t expect any characterization of DPP kernels that is as easy to check as just computing the eigenvalues of $K$ like in the symmetrical case. Instead, we give the following characterizations based on [14, Theorem 4.3.4 and Theorem 4.3.9].
Theorem 2.1.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$. $K$ is a DPP kernel if and only if one of the following properties is satisfied:
  1. (i)
    $\forall p\in\{0,1\}^{n},\leavevmode\nobreak\ \det(D(p)(I_{n}-K)+D(1_{n}-p)K)\geqslant 0.$ (2.1)
  2. (ii)
    $\forall p\in]0,1[^{n},\leavevmode\nobreak\ \det(D(p)(I_{n}-K)+D(1_{n}-p)K)>0.$ (2.2)
  3. (iii)
    For all non-zero $x\in\mathbb{R}^{n}$ there exists $i\in[n]$ such that
    $x_{i}(Kx)_{i}\geqslant 0,\leavevmode\nobreak\ |(Kx)_{i}|\leqslant|x_{i}|\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ x_{i}\neq 0.$
Proof.
The proof of (iii) is directly adapted from the method used in [14, Theorem 4.3.4] while the proof of (i) and (ii) are done differently using the interpretation of (2.1) and (2.2) in the setting of DPPs.
  1. (i)
    Direct consequence of Corollary 1.9 showing that for any $S\subset[n]$,
    $\mathbb{P}(X=S)=\det\big{(}D(\mathbbm{1}_{S^{c}})(I_{n}-K)+D(\mathbbm{1}_{S})K\big{)}.$
  2. (ii)
    We first remark that $D(p)(I_{n}-K)+D(1_{n}-p)K$ is invertible for all $p\in]0,1[^{n}$ if and only if $\det(D(p)(I_{n}-K)+D(1_{n}-p)K)>0$ for all $p\in]0,1[^{n}$. This is a consequence of the continuity of the determinant and the fact that if $p=\frac{1}{2}1_{n}$ then
    $\det(D(p)(I_{n}-K)+D(1_{n}-p)K)=\det\left(\frac{1}{2}I_{n}\right)>0.$
[…]
    $\mathbb{P}(\widetilde{X}=[n])=\mathbb{P}([n]\subset\widetilde{X})=\det\big{(}D(p)(I_{n}-K)+D(1_{n}-p)K\big{)}.$
    Now, let $S\subset[n]$ such that $\mathbb{P}(X=S)>0$. Then, by definition of $\widetilde{X}$ we have
    $\mathbb{P}(\widetilde{X}=[n])\geqslant\mathbb{P}(\widetilde{X}=[n]|X=S)\mathbb{P}(X=S)=\prod_{i\notin S}p_{i}\prod_{i\in S}(1-p_{i})\mathbb{P}(X=S)>0.$
    Therefore, $\det(D(p)(I_{n}-K)+D(1_{n}-p)K)>0$ for all $p\in]0,1[^{n}$.
  3. (iii)
    We assume that there exists $p\in]0,1[^{n}$ such that $D(p)(I_{n}-K)+D(1_{n}-p)K$ is not invertible. This means that for some non-zero $x\in\mathbb{R}^{n}$ we have
    $D(p)(x-Kx)+D(1_{n}-p)Kx=0$
[…]
    Since the function $t\mapsto 2-\frac{1}{t}$ is a bijection from ]0,1[]0,1[] 0 , 1 [ to ],1[]$\infty$,1[] - ∞ , 1 [ then we can conclude that $K$ is not a DPP kernel if and only if there exists a non-zero $x\in\mathbb{R}^{n}$ such that
    $\forall i\in[n],\leavevmode\nobreak\ \exists\mu_{i}\in]-\infty,1[\leavevmode\nobreak\ \mbox{s.t.}\leavevmode\nobreak\ x_{i}=\mu_{i}(Kx)_{i}$
    which is equivalent to
    $\forall i\in[n],\leavevmode\nobreak\ x_{i}(Kx)_{i}<0\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ |(Kx)_{i}|>|x_{i}|\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ x_{i}=0.$
    As a consequence, $K$ is a DPP kernel iff for all non-zero $x\in\mathbb{R}^{n}$ there exists $i\in[n]$ such that
    $x_{i}(Kx)_{i}\geqslant 0\leavevmode\nobreak\ ,\leavevmode\nobreak\ |(Kx)_{i}|\leqslant|x_{i}|\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ x_{i}\neq 0.\qed$
A direct consequence of these characterizations is that the set of DPP kernels is a star-shaped set centered at $\frac{1}{2}I_{n}.$
Proposition 2.2.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel. Then, for all $\lambda\in[0,1]$, $(1-\lambda)K+\lambda\left(\frac{1}{2}I_{n}\right)$ is a DPP kernel.
Proof.
Let $\widetilde{K}=(1-\lambda)K+\lambda\left(\frac{1}{2}I_{n}\right)$ and $p\in]0,1[^{n}$. By Theorem 2.1 (ii) we need to show that $D(p)(I_{n}-\widetilde{K})+D(1_{n}-p)\widetilde{K}$ is invertible to conclude that $\widetilde{K}$ is a DPP kernel. We write
$D(p)(I_{n}-\widetilde{K})+D(1_{n}-p)\widetilde{K}$ $=D(p)I_{n}+D(1_{n}-2p)\widetilde{K}$
$=D(p)I_{n}+D(1_{n}-2p)\left((1_{n}-\lambda)K+\frac{\lambda}{2}I_{n}\right)$
$=D\left(p+\frac{\lambda}{2}1_{n}-\lambda p\right)I_{n}+D(1_{n}-2p-\lambda 1_{n}+2\lambda p)K$
$=D\left(p+\frac{\lambda}{2}1_{n}-\lambda p\right)(I_{n}-K)+D\left(1_{n}-\left(p+\frac{\lambda}{2}1_{n}-\lambda p\right)\right)K$
Now, for all $i\in[n],$
$\left|p_{i}+\frac{\lambda}{2}-\lambda p_{i}-\frac{1}{2}\right|=|1-\lambda|\left|\frac{1}{2}-p_{i}\right|<\frac{1}{2}$
hence $p_{i}+\frac{\lambda}{2}-\lambda p_{i}\in]0,1[$ and thus $D(p)(I_{n}-\widetilde{K})+D(1_{n}-p)\widetilde{K}$ is invertible by Theorem 2.1 (ii), concluding the proof. ∎
2.2 Generating DPP kernels
It is known that row-diagonally dominant matrices are $P_{0}$ [14, Proposition 4.5.1] and we can get a similar result for DPP kernels.
Proposition 2.3.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ such that $K$ and $I_{n}-K$ are row diagonally dominant, meaning that the diagonal elements of $K$ are in $[0,1]$ and satisfy
$\forall i\in[n],\leavevmode\nobreak\ \min(K_{i,i},1-K_{i,i})>\sum_{j\neq i}|K_{i,j}|.$
Then $K$ is a DPP kernel.
Proof.
Let $p\in\{0,1\}^{n}$ and $\widetilde{K}=D(p)(I_{n}-K)+D(1_{n}-p)K$. Then for all $i\in[n]$,
$\widetilde{K}_{i,i}=p_{i}(1-K_{i,i})+(1-p_{i})K_{i,i}\geqslant\min(K_{i,i},1-K_{i,i})$
and for all distinct $i,j\in[n]$,
$\widetilde{K}_{i,j}=p_{i}(-K_{i,j})+(1-p_{i})K_{i,j}.$
Then,
$\sum_{j\neq i}|\widetilde{K}_{i,j}|\leqslant p_{i}\sum_{j\neq i}|K_{i,j}|+(1-p_{i})\sum_{j\neq i}|K_{i,j}|\leqslant p_{i}(1-K_{i,i})+(1-p_{i})K_{i,i}=\widetilde{K}_{i,i}.$
Using the Gershgorin circle theorem we get that that any real eigenvalue $\lambda$ of $\widetilde{K}$ is non-negative hence $\det(\widetilde{K})\geqslant 0$ and by Theorem 2.1 (i) we can conclude that $K$ is a DPP kernel. ∎
It is also known that matrices with a positive definite hermitian part are $P_{0}$ [14, Proposition 4.5.1] and adapting this result for DPP kernels yields the following proposition.
Proposition 2.4.
Let $M\in\mathcal{M}_{n}(\mathbb{R})$ such that $\|M\|_{2}\leqslant 1$. Then, $K:=\frac{1}{2}(I_{n}-M)$ is a DPP kernel.
Proof.
We can write for all $x\in\mathbb{R}^{n}$,
$\langle Kx,(I_{n}-K)x\rangle=\frac{1}{4}\langle(I_{n}-M)x,(I_{n}+M)x\rangle=\frac{1}{4}\left(\|x\|^{2}-\langle x,MM^{T}x\rangle\right)\geqslant 0.$ (2.3)
This is a consequence of the largest eigenvalue of $MM^{T}$ being $\|M\|^{2}_{2}=1$ by assumption. Now, if $K$ is not a DPP kernel then by Theorem 2.1 (ii) there exists a non-zero $x\in\mathbb{R}^{n}$ and some $p\in]0,1[^{n}$ such that
$D(p)(I_{n}-K)x+D(1_{n}-p)Kx=0.$
In particular, for all $i\in[n]$ we have
$p_{i}((I_{n}-K)x)_{i}+(1-p_{i})(Kx)_{i}=0$ (2.4)
hence $(Kx)_{i}$ and $((I_{n}-K)x)_{i}$ are either both equal to 00 or of opposite sign and thus $\langle Kx,(I_{n}-K)x\rangle\leqslant 0$. Combining (2.3) and (2.4) shows that $\langle Kx,(I_{n}-K)x\rangle=0$ and by (2.4) we must have $Kx=(I_{n}-K)x=0$ and therefore $x\in\textnormal{Ker}(K)\cap\textnormal{Ker}(I_{n}-K)=\{0_{n}\}$ which contradicts the assumption that $x\neq 0_{n}$ and therefore prove the proposition. ∎
This result shows that all matrices close to $\frac{1}{2}I_{n}$ are DDP kernels. Combining this result with Proposition 2.2 illustrates that $\frac{1}{2}I_{n}$ can be seen as the center of the set of DPP kernels. It suggests that maybe it si natural to write DPP kernels as $\frac{1}{2}(I_{n}+M)$. We point out that with this notation and if $I_{n}-K$ is invertible then $L=K(I_{n}-K)^{-1}=(I_{n}+M)(I_{n}-M)^{-1}$ corresponding to the Cayley transform of $-M$.
2.3 Eigenvalues of DPP kernels
It is already well known that symmetric DPP kernels have eigenvalues in $[0,1]$. We also points out in Corollary 1.7 that the eigenvalues of a DPP characterize the distribution of its number of points which is why it is useful to understand their behaviour. First, we give some bound on these eigenvalues using the result of [15] on $P_{0}$ matrices that we recall below.
Proposition 2.5 ([15]).
Let $L\in\mathcal{M}_{n}(\mathbb{R})$ be a $P_{0}$ matrix and $\lambda\in\mathbb{C}^{*}$ be a non-zero eigenvalue of $L$. Then,
$|\arg(\lambda)|\leqslant\pi-\frac{\pi}{n}.$
In particular, any real eigenvalue of $L$ is in $\mathbb{R}_{+}$.
Adapting this result in the determinantal setting gives the following bounds on the eigenvalues of a DPP kernel.
Proposition 2.6.
Let $\lambda$ be an eigenvalue of a DPP kernel $K\in\mathcal{M}_{n}(\mathbb{R})$. Then,
$\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}+\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)\cup\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}-\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right),$ (2.5)
where we denote by $\mathcal{B}_{\mathbb{C}}(x,r)$ the complex ball centered in $x$ with radius $r$. In particular, any real eigenvalue of a DPP kernel is in $[0,1]$.
Proof.
We begin by pointing out that when $n=1$ then (2.5) becomes $\lambda\in[0,1]$ which is obvious since the only DPP kernels of size $1\times 1$ have values in $[0,1]$. We now consider that $n\geqslant 2$ and first assume that $I_{n}-K$ is invertible. In that case, $\lambda/(1-\lambda)$ is an eigenvalue of $K(I_{n}-K)^{-1}$ which is a $P_{0}$ matrix and by Proposition 2.5 we get that either $\lambda=0$ or
$\left|\arg\left(\frac{\lambda}{1-\lambda}\right)\right|\leqslant\pi-\frac{\pi}{n}.$
[…]
$\frac{\lambda}{1-\lambda}=\frac{\lambda-|\lambda|^{2}}{|1-\lambda|}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \arg\left(\frac{\lambda}{1-\lambda}\right)=\arg(\lambda-|\lambda|^{2})=\arg(a-(a^{2}+b^{2})+bi).$
Since $b>0$ then
$\arg(a-(a^{2}+b^{2})+bi)=\left\{\begin{array}[]{l}\arctan\left(\frac{b}{a-(a^{2}+b^{2})}\right)\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ a>a^{2}+b^{2}.\\ \frac{\pi}{2}\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ a=a^{2}+b^{2}.\\ \arctan\left(\frac{b}{a-(a^{2}+b^{2})}\right)+\pi\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ a<a^{2}+b^{2}.\\ \end{array}\right.$
In the case where $a\geqslant a^{2}+b^{2}$ (equivalent to $\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2},\frac{1}{2}\right)$) then $|\arg(a-(a^{2}+b^{2})+bi)|\leqslant\pi/2$ and thus $\left|\arg\left(\frac{\lambda}{1-\lambda}\right)\right|\leqslant\frac{\pi}{2}\leqslant\pi-\frac{\pi}{n}$. In the case where $a<a^{2}+b^{2}$ then
$\left|\arg\left(\frac{\lambda}{1-\lambda}\right)\right|\leqslant\pi-\frac{\pi}{n}\leavevmode\nobreak$ $\Leftrightarrow\leavevmode\nobreak\ \arctan\left(\frac{b}{a-(a^{2}+b^{2})}\right)\leqslant-\frac{\pi}{n}$
$\Leftrightarrow\leavevmode\nobreak\ \frac{b}{a-(a^{2}+b^{2})}\leqslant-\tan\left(\frac{\pi}{n}\right)$
$\Leftrightarrow\leavevmode\nobreak\ \frac{b}{\tan\left(\frac{\pi}{n}\right)}\geqslant(a^{2}+b^{2})-a$
$\Leftrightarrow\leavevmode\nobreak\ (a^{2}-a)+\left(b^{2}-\frac{b}{\tan\left(\frac{\pi}{n}\right)}\right)\leqslant 0$
$\Leftrightarrow\leavevmode\nobreak\ \left(a-\frac{1}{2}\right)^{2}+\left(b-\frac{1}{2\tan\left(\frac{\pi}{n}\right)}\right)^{2}\leqslant\frac{1}{4\sin\left(\frac{\pi}{n}\right)^{2}}.$
Now, note that $\mathcal{B}_{\mathbb{C}}(1/2,1/2)\subset\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}+\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)$ hence
$\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}+\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)$
and by symmetry over the real line we get the desired result. Finally, if $I_{n}-K$ is not invertible then there exists some $\varepsilon>0$ such that $I_{n}-pK$ is invertible for all $p\in[1-\varepsilon,1[$. Since $pK$ is a DPP kernel (corresponding to a $p$-thinning of a DPP with kernel $K$) then
$p\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}+\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)\cup\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}-\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)$
for all $p\in[1-\varepsilon,1[$ and we get the desired result by taking $\varepsilon\rightarrow 0$. ∎
We illustrate the set of possible values for the eigenvalues of $K$ and $L$ in Figure 1.
(a) Eigenvalues of $P_{0}$ matrices
(b) Eigenvalues of DPP kernels.
Figure 1: Set of eigenvalues of $P_{0}$ matrices and DPP kernels of size $n\times n$ for $n\in\{2,\cdots,7\}$.
Since the distribution of the number of points of a DPP only depends on the eigenvalues of its kernel (see Corollary 1.7) then a good method for finding the distribution of $|X|$ when $X\sim DPP(K)$ is to find a DPP kernel $\widetilde{K}$ with the same eigenvalues as $K$ but with a simpler form from which we can easily infer the distribution of its number of points. With this method we get the following two results.
Proposition 2.7.
Let $K$ be a DPP kernel with only real eigenvalues (written with multiplicities) $\lambda_{1},\cdots,\lambda_{n}\in[0,1]$. If $X\sim DPP(K)$ then $|X|$ has the same distribution as $B_{1}+\cdots+B_{n}$ where the $B_{i}$ are independent Bernoulli random variables with $B_{i}\sim b(\lambda_{i})$.
Proof.
Obviously, the matrix $D(\lambda)$ has the same eigenvalues has $K$ and is a DPP kernel corresponding to choosing if each $i\in[n]$ is in $X$ or not independently with probability $\lambda_{i}$ hence the result. ∎
Proposition 2.8.
Let $K$ be a DPP kernel with eigenvalues in $\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2},\frac{1}{2}\right)$. We write $\lambda_{1},\cdots,\lambda_{k}$ the real eigenvalues (written with multiplicities) of $K$ and $(\mu_{1},\bar{\mu}_{1}),\cdots,(\mu_{l},\bar{\mu}_{l})$ the complex pairs of eigenvalues (written with multiplicities) of $K$. If $X\sim DPP(K)$ then $|X|$ has the same distribution as $B_{1}+\cdots+B_{k}+C_{1}+\cdots+C_{l}$ where the $B_{i}$ are independent Bernoulli random variables with $B_{i}\sim b(\lambda_{i})$ and the $C_{j}$ are independent random variables in $\{0,1,2\}$ where
$\mathbb{P}(C_{j}=0)=|\mu_{j}-1|^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}(C_{j}=1)=2\left(\frac{1}{4}-\left|\mu_{j}-\frac{1}{2}\right|^{2}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}(C_{j}=2)=|\mu_{j}|^{2}.$
Proof.
We write $\mu_{j}=a_{j}+b_{j}i$ for each $j$. It can be easily seen that $K$ has the same eigenvalues as
$\widetilde{K}=\begin{pmatrix}\lambda_{1}&&0&&&0\\ &\ddots&&&&\\ 0&&\lambda_{k}&&&\\ &&&M_{1}&&0\\ &&&&\ddots&\\ 0&&&0&&M_{l}\end{pmatrix}$
where each $M_{j}$ is the $2\times 2$ matrix $\begin{pmatrix}a_{j}&-b_{j}\\ b_{j}&a_{j}\end{pmatrix}$. It is straightforward to show that each $M_{j}$ is a DPP kernel by using Proposition 3.8 or by directly computing the associated probabilities. Therefore, by Proposition 3.5 the matrix $\widetilde{K}$ is a DPP kernel and if $X\sim DPP(K)$ then $|X|$ has the same distribution has the sum of independent random variables $B_{i}\sim b(\lambda_{i})$ and $C_{j}$ where $C_{j}$ is the number of points of a DPP with kernel $M_{j}$ hence
$\left\{\begin{array}[]{l}\mathbb{P}(C_{j}=0)=\det(I_{2}-M_{j})=(a_{j}-1)^{2}+b_{j}^{2}=|\mu_{j}-1|^{2};\\ \mathbb{P}(C_{j}=2)=\det(M_{j})=a_{j}^{2}+b_{j}^{2}=|\mu_{j}|^{2};\\ \mathbb{P}(C_{j}=1)=1-|\mu_{j}|^{2}-|\mu_{j}-1|^{2}=1-\frac{|\mu_{j}-(\mu_{j}-1)|^{2}+|\mu_{j}+(\mu_{j}-1)|^{2}}{2}=\frac{1}{2}(1-|2\mu_{j}-1|^{2}).\end{array}\right.\qed$
Proposition 2.7 is a standard result for symmetric DPP kernels but shows that this results stays true for any DPP kernel with real eigenvalues. Proposition 2.8 is a direct extension of 2.7 and can be applied to any DPP kernel of the form $K=\frac{1}{2}(I_{n}-M)$ where $\|M\|_{2}\leqslant 1$ since, using the properties of the spectral norm, any eigenvalue $\lambda$ of $K$ satisfy $|2\lambda-1|\leqslant\|M\|_{2}\leqslant 1$ hence $\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2},\frac{1}{2}\right)$. These results can be extended further by using the same reasoning as in the proof of 2.8 combined with Proposition 3.3. Let $\lambda_{1},\cdots,\lambda_{n}$ bet the eigenvalues of a DPP kernel $K$ and consider a partition $S_{1},\cdots,S_{k}$ of $[n]$ such that, for each $S_{i}$, either $S_{i}$ is of size $1$ and $\lambda_{s}\in[0,1]$ where $S_{i}=\{s\}$ or $\prod_{s\in S_{i}}(X+\frac{\lambda_{s}}{\lambda_{s}-1})$ is a polynomial with non-negative coefficients. Then, the distribution of the number of points of a DPP with kernel $K$ is the same as the distribution of $C_{1}+\cdots+C_{k}$ where the $C_{i}$ are independent random variables with $C_{i}\in\{0,\cdots,|S_{i}|\}$. While such a partition always exist, finding the optimal decomposition of a polynomial in $\mathbb{R}_{+}[X]$ into product of lower order polynomials in $\mathbb{R}_{+}[X]$ is not a solved problem to our knowledge.
3 Particular cases
In this section we give some known cases or simple examples where it is easy to check whether a DPP is well defined or not.
3.1 Examples in the literature
A lot of example of determinantal point processes are detailed in [24, Section 2.5] with some of them having non-symmetrical kernels. We mention for example the Coulomb gas in [5] or the result in [4] showing that, when taking a random partition using a Plancherel measure (later generalized in [21] to Schur measures) and transforming it into a point process on $\frac{1}{2}+\mathbb{Z}$ using a natural transformation we get a DPP with a non-symmetric kernel. We also recall below the result of [3] that $1$-dependent binary processes are DPPs whose kernel can be written as an upper Hessenberg matrix.
Proposition 3.1 ([3]).
Let $X_{1},\cdots,X_{n}\in\{0,1\}$ be a $1$-dependent process $(X_{i}\perp\!\!\!\perp X_{j}$ if $|j-i|\geqslant 2$) and $X=\{i\leavevmode\nobreak\ \mbox{s.t.}\leavevmode\nobreak\ X_{i}=1\}$. Then $X$ is a DPP with kernel of the form
$K=\begin{pmatrix}*&\cdots&\cdots&*\\ -1&\ddots&&\vdots\\ &\ddots&\ddots&\vdots\\ 0&&-1&*\end{pmatrix}.$
3.2 Construction with some specific eigenvalues
It was shown in [15] how to construct some specific $P_{0}$ matrices with a given set of eigenvalues. By adapting these results into the DPP setting we can construct some DPP kernels with specific eigenvalues.
Proposition 3.2.
Let $\lambda_{1},\cdots,\lambda_{n}\in\mathbb{C}$ be the eigenvalues (taken with multiplicities) of some $P_{0}$ matrix $M$. We consider the polynomial
$P(X)=\prod_{i=1}^{n}(X+\lambda_{i})=c_{0}+c_{1}X+\cdots+c_{n-1}X^{n-1}+X^{n}$
and the kernel
$L=\begin{pmatrix}0&-1&0&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&&\ddots&\ddots&0\\ 0&\cdots&\cdots&0&-1\\ c_{0}&\cdots&\cdots&c_{n-2}&c_{n-1}\end{pmatrix}.$ (3.1)
Then $L$ is a $P_{0}$ matrix. Moreover, if $X\sim DPP_{L}(L)$ then
$\mathbb{P}(X=S)=\left\{\begin{array}[]{l}\displaystyle\frac{c_{k-1}}{c_{0}+\cdots+c_{n-1}+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S=\{k,\cdots,n\};\\ \displaystyle\frac{1}{c_{0}+\cdots+c_{n-1}+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S=\emptyset;\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{otherwise}.\end{array}\right.$
Proof.
By definition we have $c_{k}=e_{k}(\lambda_{1},\cdots,\lambda_{n})$ which is equal to the sum of the principal minor of $M$ and thus non-negative since $M$ is $P_{0}$. Moreover, notice that $-L$ is the companion matrix of $P(X)$ hence the eigenvalues of $-L$ are the root of $P(-X)$ and therefore the eigenvalues of $L$ are the $\lambda_{i}$. Now, let $S\subset[n]$. Obviously, if $n\notin S$ then $L_{S}$ is a submatrix of a triangular matrix with zero diagonal entries hence $\det(L_{S})=0$. Otherwise, we write $S=\{s_{1},\cdots,s_{k-1},s_{k}\}$ with $s_{1}<\cdots<s_{k}=n$ and we can write $L_{S}$ as
$L_{S}=\begin{pmatrix}0&&&&\\ \vdots&&M&\\ 0&&&\\ c_{s_{1}}&c_{s_{2}}&\cdots&c_{s_{k}}\end{pmatrix}.$
where $M$ is an upper triangular matrix whose diagonal elements are $L_{s_{1},s_{2}},\cdots,L_{s_{k-1},s_{k}}$ and therefore
$\det(L_{S})=(-1)^{k+1}c_{s_{1}}\det(M)=(-1)^{k+1}c_{s_{1}}\prod_{i=1}^{k-1}L_{s_{i-1},s_{i}}.$
Now, $L_{s_{i-1},s_{i}}$ is equal to $-1$ if $s_{i}=s_{i-1}+1$ and 00 otherwise hence $\det(L_{S})$ is nonzero only when $S$ is of the form $\{k,\cdots,n\}$ with $\det(L_{S})=c_{k}\geqslant 0$ in this case. We can conclude that $L$ is a $P_{0}$ matrix. ∎
As a direct consequence, computing $K=I_{n}-(I_{n}+L)^{-1}$ when $L$ is of the form (3.1) yields the following result.
Proposition 3.3.
Let $\1\lambda_{1},\cdots,\lambda_{n}\in\mathbb{C}\backslash\{1\}$ be the eigenvalues of some DPP kernel $K$. We consider the polynomial
$P(X)=\prod_{i=1}^{n}\left(X+\frac{\lambda_{i}}{1-\lambda_{i}}\right)=c_{0}+c_{1}X+\cdots+c_{n-1}X^{n-1}+X^{n}$
and the kernel
$K=\frac{1}{c_{0}+\cdots+c_{n-1}+1}\begin{pmatrix}c_{0}&c_{0}+c_{1}&\cdots&c_{0}+\cdots+c_{n-1}\\ \vdots&\vdots&&\vdots\\ c_{0}&c_{0}+c_{1}&\cdots&c_{0}+\cdots+c_{n-1}\\ \end{pmatrix}-\begin{pmatrix}0&1&\cdots&1\\ \vdots&\ddots&\ddots&\vdots\\ \vdots&&\ddots&1\\ 0&\cdots&\cdots&0\end{pmatrix}.$
Then $K$ is a DPP kernel. Moreover, if $X\sim DPP(K)$ then
$\mathbb{P}(X=S)=\left\{\begin{array}[]{l}\displaystyle\frac{c_{k-1}}{c_{0}+\cdots+c_{n-1}+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S=\{k,\cdots,n\};\\ \displaystyle\frac{1}{c_{0}+\cdots+c_{n-1}+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S=\emptyset;\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{otherwise}.\end{array}\right.$
3.3 Examples of particular couplings
We showcase in this part some particular examples of DPP couplings. We begin by recalling some already known couplings that uses symmetric kernels.
3.3.1 Some already known repulsive couplings
One of the earliest example of construction of a specific DPP coupling (using only symmetric DPP kernels) comes from [19] where the author considered a symmetric DPP kernel $K$ and the coupling with kernel
$\mathbb{K}=\begin{pmatrix}K&(K(I_{n}-K))^{1/2}\\ (K(I_{n}-K))^{1/2}&I_{n}-K\end{pmatrix}$
which has the nice property of having a deterministic number of points since $\mathbb{K}$ is a projection matrix. A more recent example by [1] consists on simulating a DPP $X$ with some symmetric kernel $K$ and then splitting each point of $X$ into two subsets $X_{1}$ and $X_{2}$ such that each point has a probability $\frac{1}{2}$ of going into each set. Such a coupling can be expressed as a DPP with the following result.
Proposition 3.4.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ such that $2K$ is a DPP kernel. Let $X\sim DPP(2K)$. We randomly split $X$ into two disjoint subsets $X_{1}$ and $X_{2}$ such that for each $i\in X$ we choose independently and with probability $1/2$ whether $i\in X_{1}$ or $i\in X_{2}$. Then, $(X_{1},X_{2})\sim DPP(\mathbb{K})$ where
[…]
Proof.
From the definition of $X_{1}$ and $X_{2}$ we have for all $S,T\subset[n]$:
$\mathbb{P}(S\subset X_{1},T\subset X_{2})=\left\{\begin{array}[]{l}0\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S\cap T\neq\emptyset;\\ \frac{1}{2^{|S\cap T|}}\mathbb{P}(S\cup T\subset X)\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S\cap T=\emptyset.\end{array}\right.$
Now, we verify that this is the case if $(X_{1},X_{2})\sim DPP(\mathbb{K})$. Obviously if $S\cap T\neq\emptyset$ then $\mathbb{K}_{S\cup(T+n)}$ has two identical rows and columns thus its determinant is equal to 00. Otherwise, if $S\cap T=\emptyset$ then
$\det(\mathbb{K}_{S\cup(T+n)})=\det(K_{S\cup T})=\frac{1}{2^{|S\cap T|}}\det(2K_{S\cup T})=\frac{1}{2^{|S\cap T|}}\mathbb{P}(S\cup T\subset X),$
concluding the proof. ∎
Note that the assumption that $K$ is symmetrical used in [1] is not necessary for such a DPP to be well-defined.
3.3.2 Complete independency
Any independent coupling of DPPs can be written as a DPP with a block triangular form.
Proposition 3.5.
Let $n_{1},\cdots,n_{k}\in\mathbb{N}^{*}$ and write $n=n_{1}+\cdots+n_{k}$. Consider matrices $K^{(i,j)}\in\mathcal{M}_{n_{i},n_{j}}(\mathbb{R})$ with $j\geqslant i$ and define the block triangular matrix
$\mathbb{K}:=\begin{pmatrix}K^{(1,1)}&\cdots&K^{(1,k)}\\ &\ddots&\vdots\\ 0&&K^{(k,k)}\end{pmatrix}\in\mathcal{M}_{n}(\mathbb{R}).$
Then, $\mathbb{K}$ is a DPP kernel if and only if $K^{(i,i)}$ is a DPP kernel for all $i\in[k]$. In this case, if we define $\Omega_{i}=\left\{1+\sum_{j=1}^{i-1}n_{j},\cdots,\sum_{j=1}^{i}n_{j}\right\}$ and $X_{i}:=X\cap\Omega_{i}$ then all $X_{i}$ are independent with $X_{i}\sim DPP(K^{(i,i)})$.
Proof.
Direct consequence of
$\mathbb{P}(S\subset X)=\det(\mathbb{K}_{S})=\prod_{i=1}^{k}\det(\mathbb{K}_{S\cap\Omega_{i}})=\prod_{i=1}^{k}\mathbb{P}(S\cap\Omega_{i}\subset X_{i}).\qed$
Since the $K^{(i,j)}$ with $j>i$ have no influence in this case then it is more common to simply write $\mathbb{K}$ as the block diagonal matrix of the $K^{(i,i)}$.
3.3.3 Maximum dependency
We can consider that the most attractive coupling possible of two DPPs with the same distribution is a coupling of $(X_{1},X_{2})$ where $X_{1}=X_{2}$ almost surely. Such a coupling has a determinantal form as a consequence of the following two results.
Proposition 3.6.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel and define
$\mathbb{K}=\begin{pmatrix}K&I_{n}-K\\ K&I_{n}-K\end{pmatrix}.$
Then $\mathbb{K}$ is a DPP kernel. If $(X_{1},X_{2})\sim DPP(\mathbb{K})$ then $X_{1}\sim DPP(K)$, $X_{2}\sim DPP(I-K)$ and
$\mathbb{P}(X_{1}=X_{2}^{c})=1.$
Proof.
Let $X_{1}\sim DPP(K)$. It suffices to show that $\mathbb{P}(S\subset X_{1},T\subset X_{1}^{c})=0=\det(\mathbb{K}_{S\cup(T+n)})$ for any $S,T\subset[n]$ in order to prove the proposition.
  • If $S\cap T\neq\emptyset$ then $\mathbb{K}_{S\cup(T+n)}$ has two identical rows and thus $\det(\mathbb{K}_{S\cup(T+n)})=0$ hence
    $\mathbb{P}(S\subset X_{1},T\subset X_{1}^{c})=0=\det(\mathbb{K}_{S\cup(T+n)}).$
  • If $S\cap T=\emptyset$ then
    $\det(\mathbb{K}_{S\cup(T+n)})=\det\begin{pmatrix}K_{S}&-K_{S,T}\\ K_{T,S}&I_{T}-K_{T}\end{pmatrix}.$
    Using the particle hole involution (Proposition 1.8) we get
    $\det(\mathbb{K}_{S\cup(T+n)})=\mathbb{P}(S\subset X_{1},T\cap X_{1}=\emptyset)=\mathbb{P}(S\subset X_{1},T\subset X_{1}^{c}).$
Proposition 3.7.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel and define
$\mathbb{K}=\begin{pmatrix}K&I-K\\ -K&K\end{pmatrix}.$
Then $\mathbb{K}$ is a DPP kernel. If $(X_{1},X_{2})\sim DPP(\mathbb{K})$ then $X_{1},X_{2}\sim DPP(K)$ and
$\mathbb{P}(X_{1}=X_{2})=1.$
Proof.
Direct consequence of applying the particle-hole transform (Proposition 1.8) on $X_{2}$ in Proposition 3.6. ∎
3.3.4 Transformation of a symmetric coupling
By using the particle-hole transform we can show a first example of general construction of a coupling between any two DPPs with symmetric kernels.
Proposition 3.8.
Let $K^{(1)}$ and $K^{(2)}$ be two $n\times n$ symmetric matrices with eigenvalues in $[0,1]$ (and thus DPP kernels). Let $N\in\mathcal{M}_{n}(\mathbb{R})$ and define the matrix
$\mathbb{K}=\begin{pmatrix}K^{(1)}&-N\\ N^{T}&K^{(2)}\end{pmatrix}.$
Then, $\mathbb{K}$ is a DPP kernel if and only if the symmetric matrix
$\mathbb{\widetilde{K}}=\begin{pmatrix}K^{(1)}&N\\ N^{T}&I_{n}-K^{(2)}\end{pmatrix}$
has eigenvalues in $[0,1]$.
Proof.
This is a direct consequence of Proposition 1.8 with $S=\{n+1,\cdots,2n\}$. ∎
In this specific case we don’t need to used the general results established in Section 2 to study these DPPs but we can simply use the usual results on DPPs with symmetric kernel combined with the fact that
$(X_{1},X_{2})\sim DPP(\mathbb{K})\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ (X_{1},X_{2}^{c})\sim DPP(\mathbb{\widetilde{K}}).$
3.4 Examples using low rank matrices
We finish this section by giving a few examples of construction of DPPs using the nice properties of low-rank matrices regarding the determinant.
3.4.1 Rank one matrix
Proposition 3.9.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be of rank $1$. We can write $K=\lambda uv^{T}$ where $\lambda$ is the only non-zero eigenvalue of $K$ and $u,v\in\mathbb{R}^{n}$ are left and right eigenvectors chosen such that $\langle u,v\rangle=1$. Then $K$ is a DPP kernel if and only if
$\forall i\in[n],\leavevmode\nobreak\ u_{i}v_{i}\geqslant 0\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \lambda\in[0,1]$
Proof.
Since $K$ is of rank one then $\det(K_{S})=0$ when $|S|\geqslant 2$. Thus, if $X$ is a DPP with kernel $K$ then $|X|\leqslant 1$ almost surely. Now, for any $i\in[n]$ we have
$\mathbb{P}(X=\{i\})=\mathbb{P}(\{i\}\subset X)=\lambda u_{i}v_{i}\leavevmode\nobreak\ \leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}(X=\emptyset)=1-\lambda\sum_{i=1}^{n}u_{i}v_{i}=1-\lambda,$
proving the result. ∎
3.4.2 Rank one perturbation of the half identity matrix
Proposition 3.10.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be of the form $K=\frac{1}{2}(I_{n}+uv^{T})$ for some vectors $u,v\in\mathbb{R}^{n}$. Then $K$ is a DPP kernel if and only if $\sum_{i=1}^{n}|u_{i}v_{i}|\leqslant 1$.
Proof.
We first assume that $\langle u,v\rangle\neq 1$. Writing $(I_{n}-K)=\frac{1}{2}(I_{n}-uv^{T})$ and using the Sherman–Morrison formula gives
$(I_{n}-K)^{-1}=2(I_{n}-uv^{T})^{-1}=2\left(I_{n}+\frac{1}{1-\langle u,v\rangle}uv^{T}\right)$
hence
$L=K(I_{n}-K)^{-1}=(I_{n}-K)^{-1}-I_{n}=I_{n}+\frac{2}{1-\langle u,v\rangle}uv^{T}.$
Now, we show that $L$ is a $P_{0}$ matrix. Taking $S\subset[n]$ we write
$\det(L_{S})=\det\left(I_{|S|}+\frac{2}{1-\langle u,v\rangle}u_{S}v^{T}_{S}\right)=\sum_{T\subset S}\det\left(\frac{2}{1-\langle u,v\rangle}u_{T}v^{T}_{T}\right).$
Since $uv^{T}$ is a rank one matrix then all its principal minors of size $k\times k$ with $k\geqslant 2$ are equal to 00. Therefore,
$\det(L_{S})=1+\sum_{i\in S}\frac{2}{1-\langle u,v\rangle}u_{i}v_{i}.$
Moreover, using the matrix determinant Lemma gives
$\frac{1}{\det(I_{n}+L)}=\det(I_{n}-K)=\frac{1}{2^{n}}(1-\langle u,v\rangle)$
thus
$\frac{\det(L_{S})}{\det(I_{n}+L)}=\frac{1}{2^{n}}\left(1-\langle u,v\rangle+2\sum_{i\in S}u_{i}v_{i}\right)=\frac{1}{2^{n}}\left(1+\sum_{i\in S}u_{i}v_{i}-\sum_{i\not\in S}u_{i}v_{i}\right).$
This proves that the smallest possible value for $\det(L_{S})$ is $2^{-n}\det(I_{n}+L)\left(1-\sum_{i=1}^{n}|u_{i}v_{i}|\right)$ which is nonnegative if and only if $\sum_{i=1}^{n}|u_{i}v_{i}|\leqslant 1$ proving the result. We get the proof for the case $\langle u,v\rangle=1$ using a limit argument. ∎
Note that the only non-zero singular value of $uv^{T}$ is $\|u\|_{2}\|v\|_{2}$ and, by the Cauchy-Schwartz inequality, $\sum_{i=1}^{n}|u_{i}v_{i}|=\|u\|_{2}\|v\|_{2}$ only when $u$ and $v$ are proportional and thus $uv^{T}$ is symmetrical. This illustrates that among the rank one perturbations of the half identity matrix (and likely all matrices) the only ones for which Proposition 2.4 is a necessary and sufficient result are the symmetrical matrices. We also note that for these specific DPPs we have the identity
$\mathbb{P}(X=S)=\frac{1}{2^{n}}\left(1+\sum_{i\in S}u_{i}v_{i}-\sum_{i\not\in S}u_{i}v_{i}\right)\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \mathbb{P}(X=S)+\mathbb{P}(X=S^{c})=\frac{1}{2^{n-1}}$
which can be used to easily simulate these DPPs.
Proposition 3.11.
Let $u,v\in\mathbb{R}^{n}$ such that $\sum_{i=1}^{n}|u_{i}v_{i}|\leqslant 1$. Let $S$ be a random subset of $[n]$ with a uniform distribution on $\mathcal{P}([n])$. We define the random subset $X$ of $[n]$ conditionally to $S$ by
$\mathbb{P}(X=T|S)=\left\{\begin{array}[]{l}\frac{1}{2}\left(1+\sum_{i\in S}u_{i}v_{i}-\sum_{i\not\in S}u_{i}v_{i}\right)\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ T=S;\\ \frac{1}{2}\left(1+\sum_{i\not\in S}u_{i}v_{i}-\sum_{i\in S}u_{i}v_{i}\right)\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ T=S^{c};\\ 0\leavevmode\nobreak\ \mbox{otherwise.}\end{array}\right.$
$X$ is then a DPP with kernel $K=\frac{1}{2}(I_{n}+uv^{T})$.
Proof.
Direct consequence of the fact that for all $T\in[n]$ we have
$\mathbb{P}(X=T)=\mathbb{P}(X=T|S=T)\mathbb{P}(S=T)+\mathbb{P}(X=T|S=T^{c})\mathbb{P}(S=T^{c})$
and $\mathbb{P}(S=T)=\mathbb{P}(S=T^{c})=\frac{1}{2^{n-1}}$
Finally, for these DPPs we can also directly express the distribution of its number of points.
Proposition 3.12.
Let $X$ be a DPP with kernel $K\in\mathcal{M}_{n}(\mathbb{R})$ of the form $K=\frac{1}{2}(I_{n}+uv^{T})$ where $u,v\in\mathbb{R}^{n}$ satisfy $\sum_{i=1}^{n}|u_{i}v_{i}|\leqslant 1$. Then
$\mathbb{P}(|X|=k)=\frac{1}{2^{n}}\left(\binom{n}{k}+\left(\binom{n-1}{k}-\binom{n-1}{k-1}\right)\langle u,v\rangle\right).$
Proof.
$\mathbb{P}(|X|=k)$ $=\sum_{\begin{subarray}{c}S\subset[n]\\ |S|=k\end{subarray}}\mathbb{P}(X=S)$
$=\frac{1}{2^{n}}\sum_{\begin{subarray}{c}S\subset[n]\\ |S|=k\end{subarray}}\left(1-\langle u,v\rangle+2\sum_{i\in S}u_{i}v_{i}\right)$
$=\frac{1}{2^{n}}\left(\binom{n}{k}(1-\langle u,v\rangle)+2\sum_{\begin{subarray}{c}S\subset[n]\\ |S|=k\end{subarray}}\sum_{i\in S}u_{i}v_{i}\right)$
$=\frac{1}{2^{n}}\left(\binom{n}{k}(1-\langle u,v\rangle)+2\sum_{i=1}^{n}u_{i}v_{i}\sum_{\begin{subarray}{c}S\subset[n]\\ |S|=k\\ i\in S\end{subarray}}1\right)$
$=\frac{1}{2^{n}}\left(\binom{n}{k}(1-\langle u,v\rangle)+2\binom{n-1}{k-1}\langle u,v\rangle\right)$
$=\frac{1}{2^{n}}\left(\binom{n}{k}+\left(\binom{n-1}{k}-\binom{n-1}{k-1}\right)\langle u,v\rangle\right).\qed$
4 Mixing of DPPs
The most common method used to simulate DPPs with a symmetric kernel $K$ is based on a first result stating that all the eigenvalues $\lambda_{i}$ of $K$ can be replaced with independent Bernoulli random variables with parameter $\lambda_{i}$ (see [13, Theorem 4.5.3] for example), turning $K$ into a projection matrix. Then, we can use a specific algorithm for simulating DPPs with a projection matrix kernel (see [13, Algorithm 4.4.2] for example). Such a method cannot be used for DPPs with a nonsymmetric kernel for various reasons. Instead, we are currently investigating a different method for kernels of the form $K=\frac{1}{2}(I_{n}+M)$ where $\|M\|_{2}\leqslant 1$. Similarly to the method used for symmetric kernels we show that we can replace the singular values of $M$ with independent Bernoulli random variables.
Proposition 4.1.
Let $M\in\mathcal{M}_{n}(\mathbb{R})$ be a matrix such that $\|M\|_{2}\leqslant 1$ and write $M=PD(\sigma)Q^{T}$ its singular value decomposition where $P,Q$ are $n\times n$ unitary matrices and $\sigma=(\sigma_{1},\cdots,\sigma_{n})\in[0,1]^{n}$ is the vector of singular values of $M$. Let $B=(B_{1},\cdots,B_{n})$ be a vector of $n$ independent Bernoulli random variables with $B_{i}\sim b(\sigma_{i})$ and $\widetilde{K}=\frac{1}{2}(I_{n}+PD(B)Q^{T})$. Note that $\widetilde{K}$ is always a DPP kernel by Proposition 2.4. If $X$ is a DPP with kernel $\widetilde{K}$ conditionally to the $B_{i}$ then $X$ is a DPP with kernel $K=\frac{1}{2}(I_{n}+M)$.
Proof.
Let $S\subset[n]$, we can write
$\det(K_{S})=\frac{1}{2^{|S|}}\det(I_{|S|}+M_{S})=\frac{1}{2^{|S|}}\sum_{T\subset S}\det(M_{T})=\frac{1}{2^{|S|}}\sum_{T\subset S}\det(P_{T,\bullet}DQ_{\bullet,T}^{T}).$
Using the Cauchy-Binet formula twice gives
$\det(K_{S})$ $=\frac{1}{2^{|S|}}\sum_{T\subset S}\sum_{\begin{subarray}{c}U,V\subset[n]\\ |U|=|V|=|T|\end{subarray}}\det(P_{T,U})\det(D_{U,V})\det(Q_{T,V})$
$=\frac{1}{2^{|S|}}\sum_{T\subset S}\sum_{\begin{subarray}{c}U\subset[n]\\ |U|=|T|\end{subarray}}\det(P_{T,U})\det(Q_{T,V})\prod_{i\in U}\sigma_{i}.$
Similarly, we have
$\det(\widetilde{K}_{S})=\frac{1}{2^{|S|}}\sum_{T\subset S}\sum_{\begin{subarray}{c}U\subset[n]\\ |U|=|T|\end{subarray}}\det(P_{T,U})\det(Q_{T,U})\prod_{i\in U}B_{i}$
hence $\mathbb{E}[\det(\widetilde{K}_{S})]=\det(K_{S})$ and thus
$\mathbb{P}(S\subset X)=\mathbb{E}[\mathbb{P}(S\subset X|B_{1},\cdots,B_{n})]=\mathbb{E}[\det(\widetilde{K}_{S})]=\det(K_{S}).\qed$
Unfortunately, the matrix $\widetilde{K}$ doesn’t have the nice properties of projection matrices as in the symmetrical case. It is a (often) low-rank perturbation of the half identity matrix. We saw in Proposition 3.11 that its associated DPP can be easily simulated when $M$ is of rank $1$ and we are currently investigating if this could be generalized.
5 Simulating attractive couplings of DPPs
Various algorithms for simulating DPPs with non-symmetric kernels have been developed in the literature. For our simple simulations we just used the basic algorithm of [23] but more complex cases would require using some improved algorithms as in [10, 17].
5.1 Construction of the couplings
We consider DPP couplings $(X_{1},X_{2})$ on $[2n]$ such that $X_{1}$ and $X_{2}$ have the same distribution. So, if $\mathbb{K}$ is the kernel of $(X_{1},X_{2})$ then it has to be of the form
$\mathbb{K}=\begin{pmatrix}K&M\\ N&K\end{pmatrix}$ (5.1)
where $K$ is the kernel of $X_{1}$ and $X_{2}$. The cross dependency between $X_{1}$ and $X_{2}$ can be observed using the quantities
$\mathbb{P}(i\in X_{1},j\in X_{2})-\mathbb{P}(i\in X_{1})\mathbb{P}(j\in X_{2})=-M_{i,j}N_{j,i}.$
In order to have positive cross correlations we thus need to assume that at least $M_{i,i}N_{i,i}\leqslant 0$ for all $i\in[n]$. One natural way of doing it is to take $M=-N^{T}$ which corresponds to the setting of Proposition 3.8. Although, the kernel of the most attractive coupling of DPPs presented in Proposition 3.7 does not have this shape and in simulations we were only able to use such kernels to generate DPP coupling with barely any attraction. Another natural way of constructing positive cross correlations is to take $M$ (resp. $N$) to be a symmetric positive (resp. negative) semidefinite matrix. By additionally assuming that $M,N$ and $K$ commute, and are therefore simultaneously diagonalizable, and then using Proposition 2.4 we can choose $M$ and $N$ such that $\mathbb{K}$ is a DPP kernel with the following result.
Proposition 5.1.
Let $K\in\mathcal{S}^{+}_{n}(\mathbb{R})$ with eigenvalues $\lambda=(\lambda_{1},\cdots,\lambda_{n})\in[0,1]^{n}$ taken with multiplicities and write $K=PD(\lambda)P^{T}$ the spectral decomposition of $K$. Let $\mu=(\mu_{1},\cdots,\mu_{n})\in\mathbb{R}_{+}^{n}$ and $\nu=(\nu_{1},\cdots,\nu_{n})\in\mathbb{R}_{+}^{n}$ such that
$(\lambda_{i}-1/2)^{2}+\frac{1}{2}\left(\mu_{i}^{2}+\nu_{i}^{2}\pm|\mu_{i}-\nu_{i}|\sqrt{4\left(\lambda_{i}-\frac{1}{2}\right)^{2}+(\mu_{i}+\nu_{i})^{2}}\right)\leqslant\frac{1}{4}$ (5.2)
and define $M=PD(\mu)P^{T}\in\mathcal{S}^{+}_{n}(\mathbb{R})$ and $-N=PD(\nu)P^{T}\in\mathcal{S}^{+}_{n}(\mathbb{R})$. Then $\mathbb{K}=\begin{pmatrix}K&M\\ N&K\end{pmatrix}$ is a DPP kernel.
Proof.
We can write $\mathbb{K}$ as
$\mathbb{K}$ $=\begin{pmatrix}P&0\\ 0&P\end{pmatrix}\begin{pmatrix}D(\lambda)&D(\mu)\\ -D(\nu)&D(\lambda)\end{pmatrix}\begin{pmatrix}P^{T}&0\\ 0&P^{T}\end{pmatrix}$
$=\frac{1}{2}I_{2n}+\mathbb{M}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{M}=\begin{pmatrix}P&0\\ 0&P\end{pmatrix}\begin{pmatrix}D(\lambda-\frac{1}{2}1_{n})&D(\mu)\\ -D(\nu)&D(\lambda-\frac{1}{2}1_{n})\end{pmatrix}\begin{pmatrix}P^{T}&0\\ 0&P^{T}\end{pmatrix}.$
We then have
$\mathbb{M}\mathbb{M}^{T}=\begin{pmatrix}P&0\\ 0&P\end{pmatrix}\begin{pmatrix}D((\lambda-\frac{1}{2}1_{n})^{2}+\mu^{2})&D((\mu-\nu)(\lambda-\frac{1}{2}1_{n}))\\ D((\mu-\nu)(\lambda-\frac{1}{2}1_{n}))&D((\lambda-\frac{1}{2}1_{n})^{2}+\nu^{2})\end{pmatrix}\begin{pmatrix}P^{T}&0\\ 0&P^{T}\end{pmatrix}$
we can then deduce that the $2n$ eigenvalues of $\mathbb{M}\mathbb{M}^{T}$ (and thus the singular values squared of $\mathbb{M}$) are the $2$ eigenvalues of the $n$ matrices
$\begin{pmatrix}(\lambda_{i}-1/2)^{2}+\mu_{i}^{2}&(\mu_{i}-\nu_{i})(\lambda_{i}-1/2)\\ (\mu_{i}-\nu_{i})(\lambda_{i}-1/2)&(\lambda_{i}-1/2)^{2}+\nu_{i}^{2}\end{pmatrix}$
for each $i\in[n]$. It is straightforward to show that they are equal to
$(\lambda_{i}-1/2)^{2}+\frac{1}{2}\left(\mu_{i}^{2}+\nu_{i}^{2}\pm|\mu_{i}-\nu_{i}|\sqrt{4\left(\lambda_{i}-\frac{1}{2}\right)^{2}+(\mu_{i}+\nu_{i})^{2}}\right).$
If these quantities are all $\leqslant 1/4$ then Proposition 2.4 shows that $\mathbb{K}$ is a DPP kernel. ∎
Note that since the first term of (5.2) has values in $[0,\frac{1}{4}]$ and the second term can be made arbitrarily small by choosing $\mu_{i}$ and $\nu_{i}$ close to 00 then it is always possible to find some $\mu$ and $\nu$ satisfying (5.2).
5.2 Numerical results
[…]
(a) Regular grid of $31\times 31$ points
(b) Subset of the random grid chosen by a DPP
Figure 2: Example of a simulation of a DPP on a $31\times 31$ regular grid of $[0,1]^{2}$ with kernel $K_{i,j}=0.02\exp\left(-\frac{\|P_{i}-P_{j}\|_{2}^{2}}{0.018}\right).$
We then show in Figures 3 and 4 two attractive couplings of two repulsive DPPs with the same kernel on a regular grid of $[0,1]^{2}$ with $31\times 31$ points. In both cases we simulate a coupling with kernel of the form (5.1) where $K_{i,j}=0.02\exp\left(-\frac{\|P_{i}-P_{j}\|_{2}^{2}}{0.018}\right)$ in Figure 3 and $K_{i,j}=0.02\left(1+\left(\frac{\|y-x\|}{0.075}\right)^{2}\right)^{-1.1}$ in Figure 4. In both cases, the matrices $M$ and $N$ were chosen using the method described in Proposition 5.1 with the $\mu_{i}$ and $\nu_{i}$ being chosen randomly among the eigenvalues satisfying (5.2).
(a) DPP coupling simulation
(b) Inclusion probability for the orange DPP conditionally to the blue DPP
Figure 3: Simulation of a coupling of two DPPs with a Gaussian Kernel.
(a) DPP coupling simulation
(b) Inclusion probability for the orange DPP conditionally to the blue DPP
Figure 4: Simulation of a coupling of two DPPs with a Cauchy Kernel. Identical points in both DPPs are shown in red.
The simulation in Figure 3 has the intended behaviour we expected with points of both DPPs being likely to be close to each other. Although, the main drawback of this method is that it is not currently possible to control the range and the strength of the attraction since it is not clear how modifying the eigenvalues of $M$ and $N$ affect the final result. This is clearly illustrated in Figure 4 which has a very strong attraction between both DPPs but which is very localized. As a result the coupling tends to have some points of the two DPPs being very close (or even identical since we are in a discrete setting) but all other points behaving almost independently from both DPPs.