### v2

# On determinantal point processes with nonsymmetric kernels

On determinantal point processes with nonsymmetric kernels
Arnaud Poinas
Laboratoire de Mathematiques et Applications, University of Poitiers, Poitiers, France
arnaud.poinas@univ-poitiers.fr
(January 15, 2025)
In $1975$, Odile Macchi [27] introduced determinantal point processes (or DPPs for short) to model fermion particles. Since then, DPPs have seen a wide range of applications not only in physics but also in random matrix theory [32] to model the distributions of eigenvalues of some classes of random matrices, in statistics [24] to model datasets of repulsive point patterns and in machine learning [22] to select a diverse subset of a large dataset, just to name a few examples. In almost all cases, the determinantal point process is defined through a symmetric kernel and the symmetry is an important assumption needed for a lot of results (well-definedness, simulation, distribution of the number of points, …). Although, the symmetry of the kernel is not always a necessary assumption as some examples of DPPs with nonsymmetric kernels are known [6, 32, 26] and there has been a few papers working with them [11, 15]. Nevertheless, these works only focus on some specific questions related to these DPPs but does not establish their general properties. Our goal in this paper is therefore to investigate how the usual results on DPPs are modified when considering a nonsymmetric kernel.
In Section 1 we recall the basic definitions of DPPs over a finite set and some of their main properties that does not require the kernel symmetry to be satisfied. In Section 2 we use the known results on $P_{0}$ matrices to get necessary and sufficient conditions for a DPP to be well-defined and to show the behavior of the eigenvalues of DPP kernels. In Section 3 we generalize to generic DPP kernels some of the common properties of DPPs that needs the kernel symmetry. Finally, in Section 4 we take a look at how we can use the results established in the previous sections to construct and simulate attractive couplings of repulsive DPPs.
1 Definitions and first properties of DPPs
1.1 Notations and definitions
For any integer $\0n\in\mathbb{N}\backslash\{0\}$ we write $[n]$ for the set $\{1,\cdots,n\}$, $\mathcal{P}([n])$ for the power set of $[n]$, $\mathcal{M}_{n}(\mathbb{R})$ for the set of $n\times n$ matrices with real entries and $\mathcal{S}_{n}(\mathbb{R})$ (resp. $\mathcal{S}^{+}_{n}(\mathbb{R})$) for the set of symmetric (resp. symmetric positive semi-definite) matrices with real entries. For any finite set $S$ we write $|S|$ for its cardinal. Given any matrix $M\in\mathcal{M}_{n}(\mathbb{R})$ we write $\|M\|_{2}$ for its spectral norm. For any $S,T\subset[n]$ we define $M_{S,T}$ as the submatrix of $M$ with rows indexed by $S$ and columns indexed by $T$. When $S=T$ we write $M_{S}$ instead of $M_{S,S}$ in order to simplify the notations. We denote by $I_{n}$ the $n\times n$ identity matrix and $1_{n}$ (resp. $0_{n}$) the vector of $\mathbb{R}^{n}$ uniquely composed of $1$s (resp. 00s). More generally, for any $S\subset[n]$ we write $\mathbbm{1}_{S}$ for the vector whose elements indexed by $S$ are equal to $1$ and the others are equal to 00. For any vector $x=(x_{1},\cdots,x_{n})^{T}\in\mathbb{R}^{n}$ we write $D(x)$ for the $n\times n$ diagonal matrix whose diagonal elements are $x_{1},\cdots,x_{n}$. Finally, we write $X\sim b(p)$ to indicate that $X$ is a random variable with a Bernoulli distribution with parameter $p\in[0,1]$.
We begin by recalling the general definition of determinantal measures and $L$-ensemble measures.
Definition 1.1.
Let $\0n\in\mathbb{N}\backslash\{0\}$.
  • Let $L\in\mathcal{M}_{n}(\mathbb{R})$ such that $I_{n}+L$ is invertible. The L-ensemble measure $\mu$ on $\mathcal{P}([n])$ with kernel $L$ is defined by
    $\forall S\subset[n],\leavevmode\nobreak\ \mu(\{S\})=\frac{\det(L_{S})}{\det(I_{n}+L)}.$
  • Let $K\in\mathcal{M}_{n}(\mathbb{R})$. The determinantal measure $\mu$ on $\mathcal{P}([n])$ with kernel $K$ is defined by
    $\forall S\subset[n],\leavevmode\nobreak\ \mu(\{X\in\mathcal{P}([n])\leavevmode\nobreak\ \mbox{s.t.}\leavevmode\nobreak\ S\subset X\})=\det(K_{S}).$
As shown in [22] these two definitions are almost equivalent since, when $I_{n}-K$ is invertible, a determinantal measure with kernel $K$ is an $L$-ensemble measure with kernel $L=K(I_{n}-K)^{-1}$. Reciprocally, an $L$-ensemble measure with kernel $L$ is a determinantal measure with kernel $K=L(I_{n}+L)^{-1}$. These two measures are always well defined and with total mass equal to $1$ as a consequence of identity [17, 1.2.P20]:
$\forall M\in\mathcal{M}_{n}(\mathbb{R}),\leavevmode\nobreak\ \sum_{S\subset[n]}\det(M_{S})=\det(I_{n}+M).$ (1.1)
Therefore, these measures are probability measures if and only if $\mu(\{S\})\geqslant 0$ for all $S\subset[n]$. The associated probability distributions are called $L$-ensemble and determinantal point processes and we write them $\textnormal{DPP}_{L}(L)$ and $\textnormal{DPP}(K)$ respectively. We also write $X\sim\textnormal{DPP}(K)$ (resp. $X\sim\textnormal{DPP}_{L}(L)$) for a determinantal point process (resp. $L$-ensemble) with kernel $K$ (resp. $L$). When $K$ and $L$ are symmetric matrices then the condition needed for the associated probability measure to be well-defined is already known (see [22] for example).
Proposition 1.2.
Let $K,L\in\mathcal{S}_{n}(\mathbb{R})$. The determinantal measure with kernel $K$ is a probability measure if and only if the eigenvalues of $K$ are all in $[0,1]$. The $L$-ensemble measure with kernel $L$ is a probability measure if and only if $L\in\mathcal{S}_{n}^{+}(\mathbb{R})$.
For general kernels $L$, it is easy to see that its associated $L$-ensemble measure is a probability distribution if and only if its principal minors are non-negative. A matrix satisfying this property is called a $P_{0}$ matrix.
Proposition 1.3.
The $L$-ensemble measure with kernel $L\in\mathcal{M}_{n}(\mathbb{R})$ is a probability measure if and only if $L$ is a $P_{0}$ matrix, meaning that
$\forall S\subset[n],\leavevmode\nobreak\ \leavevmode\nobreak\ \det(L_{S})\geqslant 0.$
Proof.
We already mentioned that the $L$-ensemble measure is a probability measure if and only if
$\forall S\subset[n],\leavevmode\nobreak\ \leavevmode\nobreak\ \mu(\{S\})=\frac{\det(L_{S})}{\det(I_{n}+L)}\geqslant 0.$ (1.2)
Taking $S=\emptyset$ shows that we need $\det(I_{n}+L)$ to be positive hence (1.2) is equivalent to $\det(L_{S})\geqslant 0$ for all $S\subset[n]$ and $\det(I_{n}+L)\geqslant 0$. Using (1.1) we can then conclude that (1.2) is simply equivalent to $L$ having its principal minors be non-negative. ∎
On the opposite, there is no already well-known matrix family corresponding to the set of matrix $K$ whose associated determinantal measure is a probability distribution. In the rest of a paper we call such matrices a DPP kernel.
1.2 Properties of DPPs not needing the kernel symmetry
We recall some of the basic properties of DPPs that can be proved without using the assumption that their kernel is symmetrical.
Proposition 1.4.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel with complex eigenvalues (taken with multiplicities) $\lambda_{1},\cdots,\lambda_{n}\in\mathbb{C}$ and define $X\sim\textnormal{DPP}(K)$. Then,
  1. (a)
    For any $S\subset[n]$, $X\cap S\sim\textnormal{DPP}(K_{S})$. In particular, $K_{S}$ is a DPP kernel.
  2. (b)
    Let $p\in[0,1]$ and let $Y$ be the $p$-thinning of $X$$Y$ is obtained by removing each point of $X$ independently with probability $p$) then $Y\sim\textnormal{DPP}(pK)$ and thus $pK$ is a DPP kernel.
  3. (c)
    $X^{c}\sim\textnormal{DPP}(I_{n}-K)$. In particular, $I_{n}-K$ is a DPP kernel.
  4. (d)
    If $K$ is a block diagonal matrix with diagonal blocks $K_{1},\cdots,K_{l}$ then $K$ being a DPP kernel is equivalent to each $K_{i}$ being a DPP kernel.
  5. (e)
    $\mathbb{E}\left[\binom{|X|}{k}\right]=e_{k}(\lambda_{1},\cdots,\lambda_{n}),$
    where $e_{k}$ is the $k$-th elementary symmetric polynomial.
  6. (f)
    For any $x=(x_{1},\cdots,x_{n})\in\mathbb{R}^{n}$,
    $\mathbb{E}\left[\prod_{i\in X}x_{i}\right]=\det(I_{n}-D(1_{n}-x)K).$
  7. (g)
    For any $i\in[n]$ such that $K_{i,i}\neq 0$, the distribution of $X\backslash\{i\}$ conditionally to $i\in X$ is a DPP with kernel
    $\widetilde{K}=K-\frac{1}{K_{i,i}}K_{\bullet,i}K_{i,\bullet}.$ (1.3)
    In particular, $\widetilde{K}$ is a DPP kernel.
Proof.
(a) is an immediate consequence of the definition of DPPs. (b) is the direct result of $\det((pK)_{S})=p^{|S|}\det(K_{S})$. (c) comes from identity 1.1 applied to $K_{S}$:
$\det((I_{n}-K)_{S})=\sum_{T\subset S}(-1)^{|T|}\det(K_{T})$
combined with the inclusion-exclusion principle. (d) comes directly from the fact that the determinant of a block diagonal matrix is the product of the determinants of its blocks. For (e), note that $\binom{|X|}{k}$ is the number of subsets of $[n]$ with size $k$ thus
$\mathbb{E}\left[\binom{|X|}{k}\right]=\sum_{\begin{subarray}{c}S\subset[n]\\ |S|=k\end{subarray}}\mathbb{P}(S\subset X)=\sum_{\begin{subarray}{c}S\subset[n]\\ |S|=k\end{subarray}}\det(K_{S}).$
It is well-known [17, Theorem 1.2.16] that the sum of principal minors of size $k$ of any matrix $K$ is equal to $e_{k}(\lambda_{1},\cdots,\lambda_{n})$ hence the result. (f) is a particular case of the Laplace transform of DPPs shown in [31] and can be proved using (1.1) and the decomposition
$\prod_{i\in X}x_{i}=\prod_{i\in X}(1-(1-x_{i}))=\sum_{S\subset[n]}(-1)^{|S|}\left(\prod_{i\in S}(1-x_{i})\right)\mathbbm{1}_{S\subset X}.$
Finally, 1.3 is a direct application of the Schur complement (see [31, Theorem 6.5] for example). ∎
1.3 The particle-hole involution and the principal pivot transform
One of the earliest known way to create DPPs with nonsymmetric kernels is to use what is often called the particle-hole transformation. For any subset $S$ of $[n]$, the particle-hole transformation with respect to a set $S\subset[n]$ is the involution $X\mapsto(X\cap S^{c})\cup(X^{c}\cap S)$ that switches the states of the points in and out of $X\cap S$. It is shown in [7] that determinantal distributions are stable by this operation.
Proposition 1.5 ([7]).
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel and let $X\sim\textnormal{DPP}(K)$. For a given set $S\subset[n]$ we define
$\widetilde{X}=(X\cap S^{c})\cup(X^{c}\cap S).$
Then, $\widetilde{X}\sim\textnormal{DPP}(\widetilde{K})$ with
$\widetilde{K}=D(\mathbbm{1}_{S})(I_{n}-K)+D(\mathbbm{1}_{S^{c}})K.$ (1.4)
After some permutations of the rows and columns of $K$ and $\widetilde{K}$ we can write
$K=\begin{pmatrix}K_{S}&K_{S,S^{c}}\\ K_{S^{c},S}&K_{S^{c}}\end{pmatrix}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \widetilde{K}=\begin{pmatrix}I_{|S|}-K_{S}&-K_{S,S^{c}}\\ K_{S^{c},S}&K_{S^{c}}\end{pmatrix}.$ (1.5)
As a direct consequence we get an expression of all probabilities $\mathbb{P}(X=S)$ using only the kernel $K$:
Corollary 1.6 (Identity (147) in [22]).
$\mathbb{P}(X=S^{c})=\mathbb{P}([n]\subset\widetilde{X})=\det(D(\mathbbm{1}_{S})(I_{n}-K)+D(\mathbbm{1}_{S^{c}})K)=\det\begin{pmatrix}I_{|S^{c}|}-K_{S^{c}}&-K_{S^{c},S}\\ K_{S,S^{c}}&K_{S}\end{pmatrix}.$
We now give a generalization of Proposition 1.5 where, instead of switching the state of a given set of points, we choose whether we switch the state of each point in $[n]$ or not independently with its own probability.
Proposition 1.7.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel and let $X\sim\textnormal{DPP}(K)$. Let $p_{1},\cdots,p_{n}\in[0,1]$ and $B_{i}\sim b(p_{i})$ be independent random variables. We define $\widetilde{X}$ as
$\widetilde{X}=\{i\in[n]\leavevmode\nobreak\ \mbox{s.t.}\leavevmode\nobreak\ i\in X\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B_{i}=0\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ i\notin X\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B_{i}=1\}.$
Then,
$X\sim\textnormal{DPP}(D(p)(I_{n}-K)+D(1_{n}-p)K)$
Proof.
For any $S\subset[n]$ we write
$\mathbb{P}(S\subset\widetilde{X})$ $=\sum_{T\subset S}\mathbb{P}(S\subset\widetilde{X}|B_{i}=1\leavevmode\nobreak\ \mbox{when}\leavevmode\nobreak\ i\in T\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B_{i}=0\leavevmode\nobreak\ \mbox{when}\leavevmode\nobreak\ i\in T\backslash S)\prod_{i\in T}p_{i}\prod_{i\in S\backslash T}(1-p_{i})$
$=\sum_{T\subset S}\mathbb{P}(X\cap S=S\backslash T)\prod_{i\in T}p_{i}\prod_{i\in S\backslash T}(1-p_{i}).$
Using Corollary 1.6 we get
$\mathbb{P}(S\subset\widetilde{X})$ $=\det\begin{pmatrix}I_{|T|}-K_{T}&-K_{T,S\backslash T}\\ K_{S\backslash T,T}&K_{S\backslash T}\end{pmatrix}\prod_{i\in T}p_{i}\prod_{i\in S\backslash T}(1-p_{i})$
$=\sum_{T\subset S}\det\left(\begin{pmatrix}D(p)_{T}&0\\ 0&D(1_{n}-p)_{S\backslash T}\end{pmatrix}\begin{pmatrix}I_{|T|}-K_{T}&-K_{T,S\backslash T}\\ K_{S\backslash T,T}&K_{S\backslash T}\end{pmatrix}\right)$
Since $D(p)$ and $D(1_{n}-p)$ are diagonal then the matrix inside the determinant in the last identity has its rows indexed by the elements of $T$ corresponding to the rows of $(D(p)(I_{n}-K))_{S}$ and its rows indexed by the elements of $S\backslash T$ corresponding to the rows of $(D(1_{n}-p)(K))_{S}$. Therefore, by the multilinearity of the determinant with respect to its rows we get that
$\mathbb{P}(S\subset\widetilde{X})=\det\big{(}(D(p)(I_{n}-K)+D(1_{n}-p)K)_{S}\big{)},$
concluding the proof. ∎
Note that we recover Proposition 1.5 by taking $p=\mathbbm{1}_{S}$ for some $S\subset[n]$. We now show that the particle hole involution is closely linked to a common tool appearing in the theory of $P_{0}$ matrices called the principal pivot transform [33].
Definition 1.8.
Let $M\in\mathcal{M}_{n}(\mathbb{C})$ and $S\subset[n]$ such that $M_{S}$ is invertible. With the right permutation of rows and columns we can write $M$ as $\begin{pmatrix}M_{S}&M_{S,S^{c}}\\ M_{S^{c},S}&M_{S^{c}}\end{pmatrix}.$ The principal pivot transform of $M$ relative to $S$ is then defined as
$ppt(M,S):=\begin{pmatrix}M_{S}^{-1}&-M_{S}^{-1}M_{S,S^{c}}\\ M_{S^{c},S}M_{S}^{-1}&M_{S^{c}}-M_{S^{c},S}M_{S}^{-1}M_{S,S^{c}}\end{pmatrix}.$
The definition of the principal pivot transform is often better understood through the following result.
Proposition 1.9.
Let $M\in\mathcal{M}_{n}(\mathbb{C})$ and $S\subset[n]$ such that $M_{S}$ is invertible and $x,y\in\mathbb{R}^{n}$. We use the same permutation as in Definition 1.8 ans write $x=\begin{pmatrix}x_{S}\\ x_{S^{c}}\end{pmatrix}$ and $y=\begin{pmatrix}y_{S}\\ y_{S^{c}}\end{pmatrix}$. Then,
$\begin{pmatrix}y_{S}\\ y_{S^{c}}\end{pmatrix}=M\begin{pmatrix}x_{S}\\ x_{S^{c}}\end{pmatrix}\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \begin{pmatrix}x_{S}\\ y_{S^{c}}\end{pmatrix}=ppt(M,S)\begin{pmatrix}y_{S}\\ x_{S^{c}}\end{pmatrix}.$
We now show the link between the principal pivot transform and DPPs through the following result.
Proposition 1.10.
Let $L\in\mathcal{M}_{n}(\mathbb{R})$ be a $P_{0}$ matrix such that $L_{S}$ is invertible and let $X\sim DPP_{L}(L)$. For a given set $S\subset[n]$ we define the particle-hole transformation of $X$ as
$\widetilde{X}=(X\cap S^{c})\cup(X^{c}\cap S).$
If $I_{n}-\widetilde{K}$ is invertible then $\widetilde{X}\sim\textnormal{DPP}_{L}(ppt(L,S))$.
Proof.
Let $K=L(I_{n}+L)^{-1}$ be the DPP kernel of $X$ and $\widetilde{K}$ be the DPP kernel of $\widetilde{X}$. After some row and column permutations we write $K$ and $\widetilde{K}$ as in identity (1.5). Since $I_{n}-\widetilde{K}$ is invertible we can define $\widetilde{L}=(I_{n}-\widetilde{K})^{-1}-I_{n}$. We then choose $x,y\in\mathbb{R}^{n}$ indexed as in Proposition 1.9 such that $y=\widetilde{L}x$. Then,
$y=\widetilde{L}x\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ (x+y)=(I_{n}+\widetilde{L})x\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ (I_{n}-\widetilde{K})(x+y)=x\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ y=\widetilde{K}(x+y).$
This expression can be extended into
$\begin{pmatrix}y_{S}\\ y_{S^{c}}\end{pmatrix}=\begin{pmatrix}x_{S}+y_{S}-K_{S}(x_{S}+y_{S})-K_{S,S^{c}}(x_{S^{c}}+y_{S^{c}})\\ K_{S^{c},S}(x_{S}+y_{S})+K_{S^{c}}(x_{S^{c}}+y_{S^{c}})\end{pmatrix}$
[…]
We conclude that $\widetilde{L}=ppt(L,S)$ with Proposition 1.9.
The result that determinantal measures are stable by particle-hole transformations [7] is thus almost equivalent to the result that $P_{0}$ matrices are stable by principal pivot transforms [20, Theorem 4.8.4]. Interestingly, both results were proven almost at the same time but using completely different methods. This is a nice illustration of the close link there is between the theory of $P_{0}$ matrices and the theory of DPPs.
2 Properties of DPPs with nonsymmetric kernel
2.1 Characterization of DPP kernels
The first hurdle in working with nonsymmetric kernels is the issue of verifying whether a given matrix is a DPP kernel or not. In the case of $L$-ensemble, there is already a lot of literature on the problem of testing if a matrix is $P_{0}$ or not and we refer to [20, Section 4.6] for an overview. We focus instead on adapting these results for $P_{0}$ matrices to characterize whether a given matrix $K$ is a DPP kernel or not. We first recall that it was proven in [10] that the problem of checking if a given matrix is a $P$ matrix (matrix with positive principal minors) is co-NP-complete. The same can be expected for checking if a given matrix is a DPP kernel since we can go from $K$ to $L$ in polynomial time. We thus can’t expect any characterization of DPP kernels that is as easy to check as just computing the eigenvalues of $K$ like in the symmetrical case. Instead, we give the following characterizations based on [20, Theorem 4.3.4 and Theorem 4.3.9].
Theorem 2.1.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$. $K$ is a DPP kernel if and only if one of the following properties is satisfied:
  1. (a)
    $\forall p\in\{0,1\}^{n},\leavevmode\nobreak\ \det(D(p)(I_{n}-K)+D(1_{n}-p)K)\geqslant 0.$ (2.1)
  2. (b)
    $\forall p\in]0,1[^{n},\leavevmode\nobreak\ \det(D(p)(I_{n}-K)+D(1_{n}-p)K)>0.$ (2.2)
  3. (c)
    For all non-zero $x\in\mathbb{R}^{n}$ there exists $i\in[n]$ such that
    $x_{i}(Kx)_{i}\geqslant 0,\leavevmode\nobreak\ |(Kx)_{i}|\leqslant|x_{i}|\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ x_{i}\neq 0.$
Proof.
The proof of (c) is directly adapted from the method used in [20, Theorem 4.3.4] while the proof of (a) and (b) are done differently using the interpretation of (2.1) and (2.2) in the setting of DPPs.
  1. (a)
    Direct consequence of Corollary 1.6.
  2. (b)
    We first remark that $D(p)(I_{n}-K)+D(1_{n}-p)K$ is invertible for all $p\in]0,1[^{n}$ if and only if $\det(D(p)(I_{n}-K)+D(1_{n}-p)K)>0$ for all $p\in]0,1[^{n}$. This is a consequence of the continuity of the determinant and the fact that if $p=\frac{1}{2}1_{n}$ then
    $\det(D(p)(I_{n}-K)+D(1_{n}-p)K)=\det\left(\frac{1}{2}I_{n}\right)>0.$
    The continuity of the determinant also shows that (b)(a)$\Rightarrow K$ is a DPP kernel. Now, let $K$ be a DPP kernel, choose $p\in]0,1[^{n}$, define $X\sim\textnormal{DPP}(K)$ and construct $\widetilde{X}$ as in Proposition 1.7. $\widetilde{X}$ is then a determinantal point process with kernel $D(p)(I_{n}-K)+D(1_{n}-p)K$. In particular, we have
    $\mathbb{P}(\widetilde{X}=[n])=\mathbb{P}([n]\subset\widetilde{X})=\det\big{(}D(p)(I_{n}-K)+D(1_{n}-p)K\big{)}.$
    Now, let $S\subset[n]$ such that $\mathbb{P}(X=S)>0$. Then, by definition of $\widetilde{X}$ we have
    $\mathbb{P}(\widetilde{X}=[n])\geqslant\mathbb{P}(\widetilde{X}=[n]|X=S)\mathbb{P}(X=S)=\prod_{i\notin S}p_{i}\prod_{i\in S}(1-p_{i})\mathbb{P}(X=S)>0.$
    Therefore, $\det(D(p)(I_{n}-K)+D(1_{n}-p)K)>0$ for all $p\in]0,1[^{n}$.
  3. (c)
    We assume that there exists $p\in]0,1[^{n}$ such that $D(p)(I_{n}-K)+D(1_{n}-p)K$ is not invertible. This means that for some non-zero $x\in\mathbb{R}^{n}$ we have
    $D(p)(I_{n}-K)x+D(1_{n}-p)Kx=0$
[…]
    Since the function $t\mapsto 2-\frac{1}{t}$ is a bijection from ]0,1[]0,1[] 0 , 1 [ to ],1[]$\infty$,1[] - ∞ , 1 [ then we can conclude that $K$ is not a DPP kernel if and only if there exists a non-zero $x\in\mathbb{R}^{n}$ such that
    $\forall i\in[n],\leavevmode\nobreak\ \exists\mu_{i}\in]-\infty,1[\leavevmode\nobreak\ \mbox{s.t.}\leavevmode\nobreak\ x_{i}=\mu_{i}(Kx)_{i}$
    which is equivalent to
    $\forall i\in[n],\leavevmode\nobreak\ x_{i}(Kx)_{i}<0\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ |(Kx)_{i}|>|x_{i}|\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ x_{i}=0.$
    As a consequence, $K$ is a DPP kernel iff for all non-zero $x\in\mathbb{R}^{n}$ there exists $i\in[n]$ such that
    $x_{i}(Kx)_{i}\geqslant 0\leavevmode\nobreak\ ,\leavevmode\nobreak\ |(Kx)_{i}|\leqslant|x_{i}|\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ x_{i}\neq 0.\qed$
A direct consequence of these characterizations is that the set of DPP kernels is a star-shaped set centered at $\frac{1}{2}I_{n}.$
Proposition 2.2.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel. Then, for all $\lambda\in[0,1]$, $(1-\lambda)K+\lambda\left(\frac{1}{2}I_{n}\right)$ is a DPP kernel.
Proof.
Let $\widetilde{K}=(1-\lambda)K+\lambda\left(\frac{1}{2}I_{n}\right)$ and $p\in]0,1[^{n}$. By Theorem 2.1 (b) we need to show that $D(p)(I_{n}-\widetilde{K})+D(1_{n}-p)\widetilde{K}$ is invertible to conclude that $\widetilde{K}$ is a DPP kernel. We write
$D(p)(I_{n}-\widetilde{K})+D(1_{n}-p)\widetilde{K}$ $=D(p)I_{n}+D(1_{n}-2p)\widetilde{K}$
$=D(p)I_{n}+D(1_{n}-2p)\left((1-\lambda)K+\frac{\lambda}{2}I_{n}\right)$
$=D\left(p+\frac{\lambda}{2}1_{n}-\lambda p\right)I_{n}+D(1_{n}-2p-\lambda 1_{n}+2\lambda p)K$
$=D\left(p+\frac{\lambda}{2}1_{n}-\lambda p\right)(I_{n}-K)+D\left(1_{n}-\left(p+\frac{\lambda}{2}1_{n}-\lambda p\right)\right)K$
Now, for all $i\in[n],$
$\left|p_{i}+\frac{\lambda}{2}-\lambda p_{i}-\frac{1}{2}\right|=|1-\lambda|\left|\frac{1}{2}-p_{i}\right|<\frac{1}{2}$
hence $p_{i}+\frac{\lambda}{2}-\lambda p_{i}\in]0,1[$ and thus $D(p)(I_{n}-\widetilde{K})+D(1_{n}-p)\widetilde{K}$ is invertible by Theorem 2.1 (b), concluding the proof. ∎
2.2 Generating DPP kernels
It is known that row-diagonally dominant matrices are $P_{0}$ [20, Proposition 4.5.1] and we can get a similar result for DPP kernels.
Proposition 2.3.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ such that $K$ and $I_{n}-K$ are row diagonally dominant, meaning that the diagonal elements of $K$ are in $[0,1]$ and satisfy
$\forall i\in[n],\leavevmode\nobreak\ \min(K_{i,i},1-K_{i,i})\geqslant\sum_{j\neq i}|K_{i,j}|.$
Then $K$ is a DPP kernel.
Proof.
Let $p\in\{0,1\}^{n}$ and $\widetilde{K}=D(p)(I_{n}-K)+D(1_{n}-p)K$. For all distinct $i,j\in[n]$ we have,
$\widetilde{K}_{i,j}=p_{i}(-K_{i,j})+(1-p_{i})K_{i,j}.$
Then,
$\sum_{j\neq i}|\widetilde{K}_{i,j}|\leqslant p_{i}\sum_{j\neq i}|K_{i,j}|+(1-p_{i})\sum_{j\neq i}|K_{i,j}|\leqslant p_{i}(1-K_{i,i})+(1-p_{i})K_{i,i}=\widetilde{K}_{i,i}.$
Using the Gershgorin circle theorem we get that that any real eigenvalue $\lambda$ of $\widetilde{K}$ is non-negative hence $\det(\widetilde{K})\geqslant 0$ and by Theorem 2.1 (a) we can conclude that $K$ is a DPP kernel. ∎
It is also known that matrices $L\in\mathcal{M}_{n}(\mathbb{R})$ satisfying $\langle x,Lx\rangle\geqslant 0$ for all $x\in\mathbb{R}^{n}$ are $P_{0}$ [20, Proposition 4.5.1] and adapting this result for DPP kernels yields the following proposition.
Proposition 2.4.
Let $M\in\mathcal{M}_{n}(\mathbb{R})$ such that $\|M\|_{2}\leqslant 1$. Then, $K:=\frac{1}{2}(I_{n}-M)$ is a DPP kernel.
Proof.
We can write for all $x\in\mathbb{R}^{n}$,
$\langle Kx,(I_{n}-K)x\rangle=\frac{1}{4}\langle(I_{n}-M)x,(I_{n}+M)x\rangle=\frac{1}{4}\left(\|x\|^{2}-\langle x,M^{T}Mx\rangle\right)\geqslant 0.$ (2.3)
This is a consequence of the largest eigenvalue of $M^{T}M$ being $\|M\|^{2}_{2}=1$ from our assumptions. Now, if $K$ is not a DPP kernel then by Theorem 2.1 (b) there exists a non-zero $x\in\mathbb{R}^{n}$ and some $p\in]0,1[^{n}$ such that
$D(p)(I_{n}-K)x+D(1_{n}-p)Kx=0.$
In particular, for all $i\in[n]$ we have
$p_{i}((I_{n}-K)x)_{i}+(1-p_{i})(Kx)_{i}=0$ (2.4)
hence $(Kx)_{i}$ and $((I_{n}-K)x)_{i}$ are either both equal to 00 or of opposite sign and thus $\langle Kx,(I_{n}-K)x\rangle\leqslant 0$. Identity (2.3) then gives $\langle Kx,(I_{n}-K)x\rangle=0$ and by (2.4) we must have $Kx=(I_{n}-K)x=0$ and thus $x\in\textnormal{Ker}(K)\cap\textnormal{Ker}(I_{n}-K)=\{0_{n}\}$ which contradicts the assumption that $x\neq 0_{n}$ and therefore proves the proposition. ∎
This result shows that all matrices close to $\frac{1}{2}I_{n}$ are DDP kernels. Propositions 2.4 and 2.2 both illustrates that $\frac{1}{2}I_{n}$ appears as the center of the set of DPP kernels and suggests that it is natural to write DPP kernels as $\frac{1}{2}(I_{n}-M)$. We later show in section 3 that the matrix $M$ satisfies a lot of interesting properties.
2.3 Eigenvalues of DPP kernels
The eigenvalues of DPP kernels are an import tool to study DPPs. We first recall that the well-definedness of DPPs with symmetric kernels is characterized by the eigenvalues being in $[0,1]$. Moreover, as a consequence of Proposition 1.4 (e), we know that all moments of the number of points of a DPP only depends on the eigenvalues of its kernel. It means that the eigenvalues of a DPP kernel $K$ (or $L$) fully characterize the distribution of the cardinal of the associated DPP. For these reasons, we focus in this section on studying the behaviour of the eigenvalues of generic DPP kernels. We begin by giving some bound on these eigenvalues using the result of [21] on $P_{0}$ matrices we recall below.
Proposition 2.5 ([21]).
Let $L\in\mathcal{M}_{n}(\mathbb{R})$ be a $P_{0}$ matrix and $\lambda\in\mathbb{C}^{*}$ be a non-zero eigenvalue of $L$. Then,
$|\arg(\lambda)|\leqslant\pi-\frac{\pi}{n}.$
In particular, any real eigenvalue of $L$ is in $\mathbb{R}_{+}$.
Adapting this result in the determinantal setting gives the following bounds on the eigenvalues of a DPP kernel.
Proposition 2.6.
Let $\lambda$ be an eigenvalue of a DPP kernel $K\in\mathcal{M}_{n}(\mathbb{R})$. Then,
$\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}+\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)\cup\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}-\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right),$ (2.5)
where we denote by $\mathcal{B}_{\mathbb{C}}(x,r)$ the complex ball centered in $x$ with radius $r$. In particular, any real eigenvalue of a DPP kernel lies in $[0,1]$.
Proof.
We begin by pointing out that when $n=1$ then (2.5) corresponds to $\lambda\in[0,1]$ which is obvious since the only DPP kernels of size $1\times 1$ have values in $[0,1]$. We now consider that $n\geqslant 2$ and first assume that $I_{n}-K$ is invertible. In that case, $\lambda/(1-\lambda)$ is an eigenvalue of $K(I_{n}-K)^{-1}$ which is a $P_{0}$ matrix and by Proposition 2.5 we get that either $\lambda=0$ or
$\left|\arg\left(\frac{\lambda}{1-\lambda}\right)\right|\leqslant\pi-\frac{\pi}{n}.$ (2.6)
[…]
$\frac{\lambda}{1-\lambda}=\frac{\lambda-|\lambda|^{2}}{|1-\lambda|}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \arg\left(\frac{\lambda}{1-\lambda}\right)=\arg(\lambda-|\lambda|^{2})=\arg(a-(a^{2}+b^{2})+bi).$
Since $b>0$ then
$\arg(a-(a^{2}+b^{2})+bi)=\left\{\begin{array}[]{l}\arctan\left(\frac{b}{a-(a^{2}+b^{2})}\right)\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ a>a^{2}+b^{2}.\\ \frac{\pi}{2}\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ a=a^{2}+b^{2}.\\ \arctan\left(\frac{b}{a-(a^{2}+b^{2})}\right)+\pi\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ a<a^{2}+b^{2}.\\ \end{array}\right.$
If $a\geqslant a^{2}+b^{2}$ (equivalent to $\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2},\frac{1}{2}\right)$) then $|\arg(a-(a^{2}+b^{2})+bi)|\leqslant\pi/2$ and thus $\left|\arg\left(\frac{\lambda}{1-\lambda}\right)\right|\leqslant\frac{\pi}{2}\leqslant\pi-\frac{\pi}{n}$. If $a<a^{2}+b^{2}$ then
$\left|\arg\left(\frac{\lambda}{1-\lambda}\right)\right|\leqslant\pi-\frac{\pi}{n}\leavevmode\nobreak$ $\Leftrightarrow\leavevmode\nobreak\ \arctan\left(\frac{b}{a-(a^{2}+b^{2})}\right)\leqslant-\frac{\pi}{n}$
$\Leftrightarrow\leavevmode\nobreak\ \frac{b}{a-(a^{2}+b^{2})}\leqslant-\tan\left(\frac{\pi}{n}\right)$
$\Leftrightarrow\leavevmode\nobreak\ \frac{b}{\tan\left(\frac{\pi}{n}\right)}\geqslant(a^{2}+b^{2})-a$
$\Leftrightarrow\leavevmode\nobreak\ (a^{2}-a)+\left(b^{2}-\frac{b}{\tan\left(\frac{\pi}{n}\right)}\right)\leqslant 0$
$\Leftrightarrow\leavevmode\nobreak\ \left(a-\frac{1}{2}\right)^{2}+\left(b-\frac{1}{2\tan\left(\frac{\pi}{n}\right)}\right)^{2}\leqslant\frac{1}{4\sin\left(\frac{\pi}{n}\right)^{2}}.$
Now, note that $\mathcal{B}_{\mathbb{C}}(1/2,1/2)\subset\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}+\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)$ hence
$\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}+\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right).$
Since $\bar{\lambda}$ also satisfies (2.6) then with the same reasoning we get the desired result. Finally, if $I_{n}-K$ is not invertible then there exists some $\varepsilon>0$ such that $I_{n}-pK$ is invertible for all $p\in[1-\varepsilon,1[$. Since $pK$ is a DPP kernel by Proposition 1.4 (b) then
$p\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}+\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)\cup\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2}-\frac{1}{2\tan\left(\frac{\pi}{n}\right)}i,\leavevmode\nobreak\ \frac{1}{2\sin(\frac{\pi}{n})}\right)$
for all $p\in[1-\varepsilon,1[$ and we get the desired result by taking $\varepsilon\rightarrow 0$. ∎
We illustrate the set of possible values for the eigenvalues of $K$ and $L$ in Figure 1.
(a) Eigenvalues of $P_{0}$ matrices
(b) Eigenvalues of DPP kernels.
Figure 1: Set of possible eigenvalues of $P_{0}$ matrices and DPP kernels of size $n\times n$ for $n\in\{2,\cdots,7\}$.
Since the distribution of the number of points of a DPP only depends on the eigenvalues of its kernel then a useful method for finding the distribution of $|X|$ where $X\sim\textnormal{DPP}(K)$ is to find a DPP kernel $\widetilde{K}$ with the same eigenvalues as $K$ but with a simpler form from which we can easily infer the distribution of its number of points. With this method we get the following two results.
Proposition 2.7.
Let $K$ be a DPP kernel with only real eigenvalues (written with multiplicities) $\lambda_{1},\cdots,\lambda_{n}\in[0,1]$. If $X\sim\textnormal{DPP}(K)$ then $|X|$ has the same distribution as $B_{1}+\cdots+B_{n}$ where the $B_{i}$ are independent Bernoulli random variables with $B_{i}\sim b(\lambda_{i})$.
Proof.
The matrix $D(\lambda)$ has the same eigenvalues as $K$ and is a DPP kernel. It corresponds to choosing if each $i\in[n]$ is in $X$ or not independently with probability $\lambda_{i}$ hence the result. ∎
Proposition 2.8.
Let $K$ be a DPP kernel with eigenvalues in $\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2},\frac{1}{2}\right)$. We write $\lambda_{1},\cdots,\lambda_{k}$ the real eigenvalues (written with multiplicities) of $K$ and $(\mu_{1},\bar{\mu}_{1}),\cdots,(\mu_{l},\bar{\mu}_{l})$ the complex pairs of eigenvalues (written with multiplicities) of $K$. If $X\sim\textnormal{DPP}(K)$ then $|X|$ has the same distribution as $B_{1}+\cdots+B_{k}+C_{1}+\cdots+C_{l}$ where $B_{i}\sim b(\lambda_{i})$ and the $C_{j}$ are random variables in $\{0,1,2\}$ satisfying
$\mathbb{P}(C_{j}=0)=|\mu_{j}-1|^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}(C_{j}=1)=2\left(\frac{1}{4}-\left|\mu_{j}-\frac{1}{2}\right|^{2}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}(C_{j}=2)=|\mu_{j}|^{2}.$
All $B_{i}$ and $C_{j}$ are mutually independent from each other.
Proof.
We write $\mu_{j}=a_{j}+b_{j}i$ for each $j$ hence $K$ has the same eigenvalues as
$\widetilde{K}=\begin{pmatrix}\lambda_{1}&&0&&&0\\ &\ddots&&&&\\ 0&&\lambda_{k}&&&\\ &&&M_{1}&&0\\ &&&&\ddots&\\ 0&&&0&&M_{l}\end{pmatrix}$
where each $M_{j}$ is the $2\times 2$ matrix $\begin{pmatrix}a_{j}&-b_{j}\\ b_{j}&a_{j}\end{pmatrix}$. It is straightforward to see that each $M_{j}$ is a DPP kernel by directly computing the associated probabilities. Therefore, by Proposition 1.4 (d) the matrix $\widetilde{K}$ is a DPP kernel and if $X\sim\textnormal{DPP}(K)$ then $|X|$ has the same distribution has the sum of independent random variables $B_{i}\sim b(\lambda_{i})$ and $C_{j}$, where $C_{j}$ is the number of points of a DPP with kernel $M_{j}$ giving us
$\left\{\begin{array}[]{l}\mathbb{P}(C_{j}=0)=\det(I_{2}-M_{j})=(a_{j}-1)^{2}+b_{j}^{2}=|\mu_{j}-1|^{2};\\ \mathbb{P}(C_{j}=2)=\det(M_{j})=a_{j}^{2}+b_{j}^{2}=|\mu_{j}|^{2};\\ \mathbb{P}(C_{j}=1)=1-|\mu_{j}|^{2}-|\mu_{j}-1|^{2}=1-\frac{|\mu_{j}-(\mu_{j}-1)|^{2}+|\mu_{j}+(\mu_{j}-1)|^{2}}{2}=\frac{1}{2}(1-|2\mu_{j}-1|^{2}).\end{array}\right.\qed$
Proposition 2.7 is a standard result when $K$ is symmetric but we can see that this result stays true for any DPP kernel with real eigenvalues. Proposition 2.8 is a direct extension of 2.7 and can be applied to any DPP kernel of the form $K=\frac{1}{2}(I_{n}-M)$ where $\|M\|_{2}\leqslant 1$ since, using the properties of the spectral norm, any eigenvalue $\lambda$ of $K$ satisfy $|2\lambda-1|\leqslant\|M\|_{2}\leqslant 1$ hence $\lambda\in\mathcal{B}_{\mathbb{C}}\left(\frac{1}{2},\frac{1}{2}\right)$. These results can be extended further by using the same reasoning as in the proof of 2.8 combined with Proposition 2.10. Let $\lambda_{1},\cdots,\lambda_{n}$ be the eigenvalues of a DPP kernel $K$ and consider a partition $S_{1},\cdots,S_{k}$ of $[n]$ such that, for each $S_{i}$, $\prod_{s\in S_{i}}(X+\frac{\lambda_{s}}{\lambda_{s}-1})$ is a polynomial with non-negative coefficients. Then, the distribution of the number of points of a DPP with kernel $K$ is the same as the distribution of $C_{1}+\cdots+C_{k}$ where the $C_{i}$ are independent random variables with $C_{i}\in\{0,\cdots,|S_{i}|\}$. While such a partition always exist, finding the optimal factorization of a polynomial with non-negative coefficients into a product of lower order polynomials with non-negative coefficients is a difficult problem to our knowledge [8].
2.4 Construction with some specific eigenvalues
It was shown in [21] how to construct some specific $P_{0}$ matrices with a given set of eigenvalues. By adapting these results into the DPP setting we can also construct some DPP kernels with a given set of eigenvalues.
Proposition 2.9.
Let $\lambda_{1},\cdots,\lambda_{n}\in\mathbb{C}$ be the eigenvalues (taken with multiplicities) of some $P_{0}$ matrix $M$. We consider the polynomial
$P(X)=\prod_{i=1}^{n}(X+\lambda_{i})=c_{0}+c_{1}X+\cdots+c_{n-1}X^{n-1}+X^{n}$
and the kernel
$L=\begin{pmatrix}0&-1&0&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&&\ddots&\ddots&0\\ 0&\cdots&\cdots&0&-1\\ c_{0}&\cdots&\cdots&c_{n-2}&c_{n-1}\end{pmatrix}.$ (2.7)
Then $L$ is a $P_{0}$ matrix. Moreover, if $X\sim DPP_{L}(L)$ then
$\mathbb{P}(X=S)=\left\{\begin{array}[]{l}\displaystyle\frac{c_{k-1}}{c_{0}+\cdots+c_{n-1}+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S=\{k,\cdots,n\};\\ \displaystyle\frac{1}{c_{0}+\cdots+c_{n-1}+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S=\emptyset;\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{otherwise}.\end{array}\right.$
Proof.
Using standard results on characteristic polynomials we have $c_{k}=e_{k}(\lambda_{1},\cdots,\lambda_{n})$ which is equal to the sum of all principal minors of $L$ of size $k$ [17, Theorem 1.2.16] and is therefore non-negative since $L$ is a $P_{0}$ matrix. Moreover, notice that $-L$ is the companion matrix of $P(X)$ hence the eigenvalues of $-L$ are the root of $P(-X)$ and therefore the eigenvalues of $L$ are the $\lambda_{i}$. Now, let $S\subset[n]$. Obviously, if $n\notin S$ then $L_{S}$ is a submatrix of a triangular matrix whose diagonal entries are all zeros hence $\det(L_{S})=0$. Otherwise, we write $S=\{s_{1},\cdots,s_{k-1},s_{k}\}$ with $s_{1}<\cdots<s_{k}=n$ and we can write $L_{S}$ as
$L_{S}=\begin{pmatrix}0&&&&\\ \vdots&&M&\\ 0&&&\\ c_{s_{1}}&c_{s_{2}}&\cdots&c_{s_{k}}\end{pmatrix}.$
where $M$ is an upper triangular matrix whose diagonal elements are $L_{s_{1},s_{2}},\cdots,L_{s_{k-1},s_{k}}$ and therefore
$\det(L_{S})=(-1)^{k+1}c_{s_{1}}\det(M)=(-1)^{k+1}c_{s_{1}}\prod_{i=1}^{k-1}L_{s_{i-1},s_{i}}.$
Now, $L_{s_{i-1},s_{i}}$ is equal to $-1$ if $s_{i}=s_{i-1}+1$ and 00 otherwise hence $\det(L_{S})$ is nonzero only when $S$ is of the form $\{k,\cdots,n\}$ with $\det(L_{S})=c_{k}\geqslant 0$ in this case. We can conclude that $L$ is a $P_{0}$ matrix. ∎
As a direct consequence, computing $K=I_{n}-(I_{n}+L)^{-1}$ when $L$ is of the form (2.7) yields the following result.
Proposition 2.10.
Let $\1\lambda_{1},\cdots,\lambda_{n}\in\mathbb{C}\backslash\{1\}$ be the eigenvalues of some DPP kernel $K$. We consider the polynomial
$P(X)=\prod_{i=1}^{n}\left(X+\frac{\lambda_{i}}{1-\lambda_{i}}\right)=c_{0}+c_{1}X+\cdots+c_{n-1}X^{n-1}+X^{n}$
and the kernel
$K=\frac{1}{c_{0}+\cdots+c_{n-1}+1}\begin{pmatrix}c_{0}&c_{0}+c_{1}&\cdots&c_{0}+\cdots+c_{n-1}\\ \vdots&\vdots&&\vdots\\ c_{0}&c_{0}+c_{1}&\cdots&c_{0}+\cdots+c_{n-1}\\ \end{pmatrix}-\begin{pmatrix}0&1&\cdots&1\\ \vdots&\ddots&\ddots&\vdots\\ \vdots&&\ddots&1\\ 0&\cdots&\cdots&0\end{pmatrix}.$
Then $K$ is a DPP kernel. Moreover, if $X\sim\textnormal{DPP}(K)$ then
$\mathbb{P}(X=S)=\left\{\begin{array}[]{l}\displaystyle\frac{c_{k-1}}{c_{0}+\cdots+c_{n-1}+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S=\{k,\cdots,n\};\\ \displaystyle\frac{1}{c_{0}+\cdots+c_{n-1}+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ S=\emptyset;\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{otherwise}.\end{array}\right.$
3 Properties of general DPPs
We saw in Proposition 2.2 that $\frac{1}{2}I_{n}$ can be seen as the center of the set of DPP kernels and we showed in Proposition 2.4 that the set of DPP kernels contains the ball centered at $\frac{1}{2}I_{n}$ with radius $\frac{1}{2}$ for the spectral norm. This suggest that a natural way of constructing DPP kernels is to write them as $K=\frac{1}{2}(I_{n}-M)$. We quickly point out that the relationship between $M$ and the likelihood kernel $L$ is a standard matrix transformation called the Cayley transform whose main property is that it maps skew-symmetric matrices on special orthogonal matrices.
Proposition 3.1.
Let $K=\frac{1}{2}(I_{n}-M)$ be a DPP kernel such that $I_{n}+M$ is invertible. Then, $L$ is the Cayley transform of $M$:
$L=(I_{n}-M)(I_{n}+M)^{-1}.$
3.1 The parity kernel
The minors of the kernels $K$ and $L$ gives, respectively, the inclusion and exact probabilities of the DPP. This is why $K$ and $L$ are often called, respectively, the correlation and likelihood kernel of a DPP. In comparison, the minors of the kernel $M$ also have a nice direct interpretation as a consequence of the Laplace transform formula (see Proposition 1.4 (f)):
$X\sim DPP(K)\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \forall S\subset[n],\leavevmode\nobreak\ \mathbb{E}\left[(-1)^{|X\cap S|}\right]=\det((I_{n}-2K)_{S})=\det(M_{S})$ (3.1)
hence
$\mathbb{P}(|X\cap S|\leavevmode\nobreak\ \mbox{is even})=\frac{1}{2}(1+\det(M_{S}))\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \mathbb{P}(|X\cap S|\leavevmode\nobreak\ \mbox{is odd})=\frac{1}{2}(1-\det(M_{S})).$
The minors of the kernel $M$ gives information about the parity of the number of points of a DPP in a given subset. For this reason we call $M$ the parity kernel of the DPP. Generalizing (3.1) gives the following proposition.
Proposition 3.2.
Let $S_{1},\cdots,S_{k}$ be disjoint subsets of $[n]$. Let $R_{1},\cdots,R_{k}$ defined by
$R_{i}=(-1)^{|X\cap S_{i}|}=\left\{\begin{array}[]{l}1\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ |X\cap S_{i}|\leavevmode\nobreak\ \mbox{is even,}\\ -1\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ |X\cap S_{i}|\leavevmode\nobreak\ \mbox{is odd.}\end{array}\right.$
Then, for all $\varepsilon\in\{-1,+1\}^{k}$,
$\mathbb{P}(\forall i,\leavevmode\nobreak\ R_{i}=\varepsilon_{i})=\frac{1}{2^{k}}\sum_{T\subset[k]}\left(\prod_{i\in T}\varepsilon_{i}\right)\det(M_{\cup_{i\in T}S_{i}})$
Proof.
Using expression (3.1) and the fact that $1+\varepsilon_{i}R_{i}$ is equal to 00 if $R_{i}\neq\varepsilon_{i}$ and $2$ if $R_{i}=\varepsilon_{i}$ we get
$\mathbb{P}(\forall i,\leavevmode\nobreak\ R_{i}=\varepsilon_{i})$ $=\frac{1}{2^{k}}\mathbb{E}\left[\prod_{i=1}^{k}\left(1+\varepsilon_{i}R_{i}\right)\right]$
$=\frac{1}{2^{k}}\mathbb{E}\left[\prod_{i=1}^{k}\left(1+\varepsilon_{i}(-1)^{|X\cap S_{i}|}\right)\right]$
$=\frac{1}{2^{k}}\sum_{T\subset[k]}\mathbb{E}\left[\prod_{i\in T}\varepsilon_{i}(-1)^{|X\cap S_{i}|}\right]$
$=\frac{1}{2^{k}}\sum_{T\subset[k]}\left(\prod_{i\in T}\varepsilon_{i}\right)\mathbb{E}\left[(-1)^{|X\cap(\cup_{i\in T}S_{i})|}\right]$
$=\frac{1}{2^{k}}\sum_{T\subset[k]}\left(\prod_{i\in T}\varepsilon_{i}\right)\det(M_{\cup_{i\in T}S_{i}}).$
As a consequence, for any $S\subset[n]$ we can write
$\left\{\begin{array}[]{l}\mathbb{P}(|X\cap S|\leavevmode\nobreak\ \mbox{even and}\leavevmode\nobreak\ |X\cap S^{c}|\leavevmode\nobreak\ \mbox{even})=\frac{1}{4}\left(1+\det(M_{S})+\det(M_{S^{c}})+\det(M)\right);\\ \mathbb{P}(|X\cap S|\leavevmode\nobreak\ \mbox{even and}\leavevmode\nobreak\ |X\cap S^{c}|\leavevmode\nobreak\ \mbox{odd})=\frac{1}{4}\left(1+\det(M_{S})-\det(M_{S^{c}})-\det(M)\right);\\ \mathbb{P}(|X\cap S|\leavevmode\nobreak\ \mbox{odd and}\leavevmode\nobreak\ |X\cap S^{c}|\leavevmode\nobreak\ \mbox{even})=\frac{1}{4}\left(1-\det(M_{S})+\det(M_{S^{c}})-\det(M)\right);\\ \mathbb{P}(|X\cap S|\leavevmode\nobreak\ \mbox{odd and}\leavevmode\nobreak\ |X\cap S^{c}|\leavevmode\nobreak\ \mbox{odd})=\frac{1}{4}\left(1-\det(M_{S})-\det(M_{S^{c}})+\det(M)\right).\end{array}\right.$
Additionally, there is an almost symmetric relationship between the probabilities of a DPP and the minors of $M$ due to the Cayley involution (Proposition 3.1) mapping the kernel $M$ to the kernel $L$.
Proposition 3.3.
Let $K=\frac{1}{2}(I_{n}-M)\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel and $X\sim\textnormal{DPP}(K)$. Then, for any $S\subset[n]$:
$\mathbb{P}(X=S)=\frac{1}{2^{n}}\sum_{A\subset[n]}(-1)^{|A\cap S|}\det(M_{A});$ (3.2)
$\det(M_{S})=\sum_{A\subset[n]}(-1)^{|A\cap S|}\mathbb{P}(X=A).$ (3.3)
Proof.
Identity (3.3) is an immediate consequence of (3.1) since
$\mathbb{E}\left[(-1)^{|X\cap S|}\right]=\sum_{A\subset[n]}(-1)^{|A\cap S|}\mathbb{P}(X=S).$
In order to reverse this identity we consider an arbitrary oredring of $\mathcal{P}([n])$ and define two vectors $u,v\in\mathbb{R}^{\mathcal{P}([n])}$ by $u_{A}=\det(M_{A})$ and $v_{A}=\mathbb{P}(X=A)$ and the family $\{e(S),S\in\mathcal{P}([n])\}$ of vectors in $\mathbb{R}^{\mathcal{P}([n])}$ with $e(S)_{A}=(-1)^{|A\cap S|}$. With these notations we can rewrite (3.3) as
$u_{S}=\det(M_{S})=\langle v,e(S)\rangle$
Since $\langle e(S),e(S)\rangle=2^{n}$ for all $S\subset[n]$, if we prove that $\{e(S),S\in\mathcal{P}([n])\}$ is an orthogonal basis of $\mathbb{R}^{\mathcal{P}([n])}$ then we can conclude that for all $S\subset[n]$:
$v=\sum_{A\subset[n]}\frac{u_{A}}{\langle e(A),e(A)\rangle}e(A)=\frac{1}{2^{n}}\sum_{A\subset[n]}u_{A}e(A)\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ v_{S}=\frac{1}{2^{n}}\sum_{A\subset[n]}u_{A}e(A)_{S},$
corresponding to (3.2).
Let $S,T\subset[n]$, then
$\langle e(S),e(T)\rangle=\sum_{A\subset[n]}(-1)^{|A\cap S|+|A\cap T|}.$
We can write any set $A\subset[n]$ as the disjoint union of $4$ sets: $U\subset S\cap T$, $V\subset S\backslash S\cap T$, $W\subset T\backslash S\cap T$ and $Z\subset(S\cup T)^{c}$ yielding
$\langle e(S),e(T)\rangle$ $=\sum_{U\subset S\cap T}\sum_{V\subset S\backslash S\cap T}\sum_{W\subset T\backslash S\cap T}\sum_{Z\subset(S\cup T)^{c}}(-1)^{|U|+|V|+|U|+|W|}$
$=\sum_{U\subset S\cap T}\sum_{Z\subset(S\cup T)^{c}}\left(\sum_{V\subset S\backslash S\cap T}(-1)^{|V|}\right)\left(\sum_{W\subset T\backslash S\cap T}(-1)^{|W|}\right)$
$=\sum_{U\subset S\cap T}\sum_{Z\subset(S\cup T)^{c}}(1-1)^{|T\backslash S\cap T|}(1-1)^{|S\backslash S\cap T|}.$
If $S\neq T$ then either $S\backslash S\cap T\neq\emptyset$ or $T\backslash S\cap T\neq\emptyset$ proving that $\langle e(S),e(T)\rangle=0$ and concluding the proof.
3.2 Mixing of orthogonal DPPs
The most common method used to simulate DPPs with a symmetric kernel $K$ is based on a first result stating that all the eigenvalues $\lambda_{i}$ of $K$ can be replaced with independent Bernoulli random variables with parameter $\lambda_{i}$ (see [18, Theorem 4.5.3] for example), turning $K$ into a projection matrix. Such a DPP is called a projection DPP and its various nice properties (such as having a deterministic number of points) allows for a simple simulation algorithm (see [18, Algorithm 4.4.2] for example) that can be used both for finite and continuous DPPs. We generalize this principle for kernels of the form $K=\frac{1}{2}(I_{n}-M)$ by proving that we can replace the singular values $\sigma_{i}$ of the parity kernel $M$ by random variables with values in $\{-1,+1\}$. The kernel $M$ then become an orthogonal matrix. We call the associated determinantal distribution an orthogonal DPP.
Proposition 3.4.
Let $M\in\mathcal{M}_{n}(\mathbb{R})$ be a matrix such that $\|M\|_{2}\leqslant 1$ and write $M=PD(\sigma)Q^{T}$ its singular value decomposition where $P,Q$ are $n\times n$ orthogonal matrices and $\sigma=(\sigma_{1},\cdots,\sigma_{n})\in[0,1]^{n}$ is the vector of singular values of $M$. Let $B=(B_{1},\cdots,B_{n})$ be a vector of $n$ independent random variables taking value in $\{-1,1\}$ and satisfying $\mathbb{E}[B_{i}]=\sigma_{i}$. Define $\widetilde{K}=\frac{1}{2}(I_{n}-PD(B)Q^{T})$. Note that $\widetilde{K}$ is always a DPP kernel by Proposition 2.4. If $X$ is a DPP with kernel $\widetilde{K}$ conditionally to $B$ then $X$ is a DPP with kernel $K=\frac{1}{2}(I_{n}-M)$.
Proof.
Let $S\subset[n]$, we can write
$\det(K_{S})=\frac{1}{2^{|S|}}\det(I_{|S|}-M_{S})=\frac{1}{2^{|S|}}\sum_{T\subset S}\det(-M_{T})=\frac{1}{2^{|S|}}\sum_{T\subset S}(-1)^{|T|}\det(P_{T,[n]}D(\sigma)Q_{[n],T}^{T}).$
Using the Cauchy-Binet formula twice and the fact that $\det(D(\sigma)_{U,V})=(\prod_{i\in U}\sigma_{i})\mathbbm{1}_{U=V}$ gives
$\det(K_{S})$ $=\frac{1}{2^{|S|}}\sum_{T\subset S}(-1)^{|T|}\sum_{\begin{subarray}{c}U,V\subset[n]\\ |U|=|V|=|T|\end{subarray}}\det(P_{T,U})\det(D(\sigma)_{U,V})\det(Q_{T,V})$
$=\frac{1}{2^{|S|}}\sum_{T\subset S}(-1)^{|T|}\sum_{\begin{subarray}{c}U\subset[n]\\ |U|=|T|\end{subarray}}\det(P_{T,U})\det(Q_{T,U})\prod_{i\in U}\sigma_{i}.$
With the same reasoning we also get
$\det(\widetilde{K}_{S})=\frac{1}{2^{|S|}}\sum_{T\subset S}(-1)^{|T|}\sum_{\begin{subarray}{c}U\subset[n]\\ |U|=|T|\end{subarray}}\det(P_{T,U})\det(Q_{T,U})\prod_{i\in U}B_{i},$
hence $\mathbb{E}[\det(\widetilde{K}_{S})]=\det(K_{S})$ and thus
$\mathbb{P}(S\subset X)=\mathbb{E}[\mathbb{P}(S\subset X|B_{1},\cdots,B_{n})]=\mathbb{E}[\det(\widetilde{K}_{S})]=\det(K_{S}).\qed$
3.3 Properties of orthogonal DPPs
We point out that a slight potential source of confusion is that the name "projection DPP" refers to the correlation kernel $K$ being a projection matrix, while "orthogonal DPP" refers to the parity matrix $M$ being orthogonal. The kernel $K$ of orthogonal DPPs satisfy
$K^{T}K=KK^{T}=\frac{1}{2}(K+K^{T}),$ (3.4)
which does not corresponds to a common class of matrix to our knowledge, except that it is a particular case of normal matrices. When $K$ is symmetric, identity (3.4) becomes $K^{2}=K$ meaning that orthogonal DPPs generalize projection DPPs. Just like projection DPPs we show that orthogonal DPPs are stable by conditioning with respect to the presence of a point.
Proposition 3.5.
Let $X$ be an orthogonal DPP with associated kernel $K=\frac{1}{2}(I_{n}-M)$, $M$ being and orthogonal matrix. Let $i\in[n]$ such that $K_{i,i}\neq 0$. Then, the distribution of $X\backslash\{i\}$ conditionally to $i\in X$ is an orthogonal DPP.
Proof.
Let $\widetilde{K}$, defined as in (1.3), be the kernel of the distribution of $X\backslash\{i\}$ conditionally to $i\in X$ and define $\widetilde{M}:=I_{n}-2\widetilde{K}$. Since $K$ satisfy (3.4) then
$KK_{i,\bullet}^{T}=\frac{1}{2}(K_{\bullet,i}+K_{i,\bullet}^{T})\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ MK_{i,\bullet}^{T}=K_{i,\bullet}^{T}-(K_{\bullet,i}+K_{i,\bullet}^{T})=-K_{\bullet,i}$
and
$\langle K_{i,\bullet}^{T},K_{i,\bullet}^{T}\rangle=\frac{1}{2}(K_{i,i}+K_{i,i})=K_{i,i}.$
Hence, if we define the unitary vector $v=\frac{1}{\|K_{i,\bullet}^{T}\|}K_{i,\bullet}^{T}$ then
$\widetilde{M}=I_{n}-2K+\frac{2}{K_{i,i}}K_{\bullet,i}K_{i,\bullet}=M-\frac{2}{\|K_{i,\bullet}^{T}\|^{2}}MK_{i,\bullet}^{T}K_{i,\bullet}=M(1-2vv^{T}),$
therefore $\widetilde{M}$ is the product of an orthogonal matrix and an householder reflection and thus an orthogonal matrix itself. ∎
More generally, we show that, unlike projection DPPs, orthogonal DPPs are also stable by particle-hole involution.
Proposition 3.6.
Let $X$ be an orthogonal DPP with associated kernel $K=\frac{1}{2}(I_{n}-M)$, $M$ being and orthogonal matrix. Let $S\subset[n]$ and define $\widetilde{X}$ as in Proposition 1.5. Then, the distribution of $\widetilde{X}$ is an orthogonal DPP.
Proof.
We saw in Proposition 1.5 that $\widetilde{X}$ is a DPP with kernel $\widetilde{K}$ satisfying identity (1.4). Thus,
$\widetilde{M}$ $=I_{n}-2\widetilde{K}$
$=I_{n}-2D(\mathbbm{1}_{S})(I_{n}-K)-2(I_{n}-D(\mathbbm{1}_{S}))K$
$=I_{n}-D(\mathbbm{1}_{S})(I_{n}+M)-(I_{n}-D(\mathbbm{1}_{S}))(I_{n}-M)$
$=I_{n}-D(\mathbbm{1}_{S})-D(\mathbbm{1}_{S})M-I_{n}+M+D(\mathbbm{1}_{S})-D(\mathbbm{1}_{S})M$
$=(I_{n}-2D(\mathbbm{1}_{S}))M$
Since $D(\mathbbm{1}_{S})$ is a symmetric projection matrix then $(I_{n}-2D(\mathbbm{1}_{S}))M$ is orthogonal. ∎
Note that, as a direct consequence of the proof of Proposition 3.6, the particle involution with respect to a singleton $\{i\}$ multiply $M$ by the householder reflection with respect to the $i$-th canonical vector.
Finally, an immediate consequence of the decomposition of real orthogonal matrices gives the following decomposition of the $K$ and $L$ kernels of orthogonal DPPs.
Proposition 3.7.
Let $K$ be the kernel of an orthogonal DPP. There exists an $n\times n$ orthogonal matrix $P$ such that $PKP^{T}$ is written as
$\begin{pmatrix}R_{1}&&0&&&\\ &\ddots&&&0&\\ 0&&R_{k}&&&\\ &&&\delta_{1}&&0\\ &0&&&\ddots&\\ &&&0&&\delta_{n-2k}\end{pmatrix}$ (3.5)
where $\delta_{i}\in\{0,1\}$ and each $R_{i}$ is a $2\times 2$ matrix of the form
$R_{i}=\begin{pmatrix}\frac{1+\cos(\theta_{i})}{2}&-\frac{\sin(\theta_{i})}{2}\\ \frac{\sin(\theta_{i})}{2}&\frac{1+\cos(\theta_{i})}{2}\end{pmatrix},\leavevmode\nobreak\ \theta_{i}\in]0,\pi[.$
If all $\delta_{i}$ are equal to 00 then $L$ is a well defined skew-symmetric matrix and $PLP^{T}$ is written
$\begin{pmatrix}R^{\prime}_{1}&&0&&&\\ &\ddots&&&0&\\ 0&&R^{\prime}_{k}&&&\\ &&&0&&0\\ &0&&&\ddots&\\ &&&0&&0\end{pmatrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ R^{\prime}_{i}=\begin{pmatrix}0&-\tan\left(\frac{\pi-\theta_{i}}{2}\right)\\ \tan\left(\frac{\pi-\theta_{i}}{2}\right)&0\end{pmatrix}.$ (3.6)
We point out that a DPP with kernel $K=\begin{pmatrix}\frac{1+\cos(\theta)}{2}&-\frac{\sin(\theta)}{2}\\ \frac{\sin(\theta)}{2}&\frac{1+\cos(\theta)}{2}\end{pmatrix}$ satisfy
$\mathbb{P}(X=\emptyset)=\frac{1-\cos(\theta)}{2},\leavevmode\nobreak\ \mathbb{P}(X=\{1\})=\mathbb{P}(X=\{2\})=0\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \mathbb{P}(X=\{1,2\})=\frac{1+\cos(\theta)}{2},$
telling us that the cardinal of orthogonal DPPs with the decomposition (3.5) is of the form $l+2B$ where $l$ is the number of $\delta_{i}$ equal to $1$ and $B$ is a Poisson-Binomial distribution with probabilities $\frac{1+\cos(\theta_{1})}{2},\cdots,\frac{1+\cos(\theta_{k})}{2}$. If $k=0$ then $K$ is a projection matrix and we recover the usual result that projection DPPs have a deterministic amount of points.
3.4 Simulation of orthogonal DPPs
By Proposition (3.5), the kernel $K$ of an orthogonal DPP has a number $k$ of pairs of complex eigenvalues, a number $l$ of eingevalues equal to $1$ and a number $n-l-2k$ of eigenvalues equal to 00. When $k=0$ this corresponds to a projection DPP for whom we already know an efficient sampling algorithm [18, Algorithm 4.4.2]. When $k=1$ we can show that the corresponding orthogonal DPP is a mixture of two projection DPPs.
Proposition 3.8.
Let $P$ be a matrix with size $n\times(l+2)$ satisfying $P^{T}P=I_{l+2}$ and $K$ be the kernel of an orthogonal DPP such that
$K=PDP^{T}\leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ D=\begin{pmatrix}\alpha&\beta&0&\cdots&0\\ -\beta&\alpha&0&\cdots&0\\ 0&0&1&&0\\ \vdots&\vdots&&\ddots&\\ 0&0&0&&1\end{pmatrix}$ (3.7)
with $\alpha\in]0,1[$ and $\alpha^{2}+\beta^{2}=\alpha$. Now, define $X$ as a projection DPP with kernel $K^{(1)}=PP^{T}$ and $Y$ as a projection DPP with kernel $K^{(2)}=P_{[n],\{3,\cdots,l+2\}}P_{[n],\{3,\cdots,l+2\}}^{T}$. Let $Z$ be a random variable defined conditionnaly to $(X,Y)$ such that $Z=X$ with probability $\alpha$ and $Z=Y$ with probability $1-\alpha$ (the choice being independent from $X$ and $Y$). Then, $Z$ is a DPP with kernel $K$.
Proof.
Let $S\subset[n]$. If $|S|\geqslant l+2$ then $\det(K_{S})=0$ since $K$ is of rank $l+2$. Otherwise, the Cauchy-Binet formula gives
$\det(K_{S})=\sum_{\begin{subarray}{c}T,U\subset[l+2]\\ |T|=|U|=|S|\end{subarray}}\det(P_{S,T})\det(D_{T,U})\det(P_{S,U}).$
Due to the shape (3.7) of the matrix $D$, $\det(D_{T,U})$ vanishes when $T\cap\{3,\cdots,l+2\}\neq U\cap\{3,\cdots,l+2\}$ or $|T\cap\{1,2\}|\neq|U\cap\{1,2\}|$ thus
$\det(K_{S})=\sum_{\begin{subarray}{c}V\subset\{3,\cdots,l+2\}\\ |V|=|S|\end{subarray}}\det(P_{S,V})^{2}+\sum_{\begin{subarray}{c}V\subset\{3,\cdots,l+2\}\\ |V|=|S|-2\end{subarray}}(\alpha^{2}+\beta^{2})\det(P_{S,V\cup\{1,2\}})^{2}\\ +\sum_{\begin{subarray}{c}V\subset\{3,\cdots,l+2\}\\ |V|=|S|-1\end{subarray}}\left(\alpha\det(P_{S,V\cup\{1\}})^{2}+\alpha\det(P_{S,V\cup\{2\}})^{2}+(\beta-\beta)\det(P_{S,V\cup\{2\}})\det(P_{S,V\cup\{1\}})\right).$
Using the identity $\alpha^{2}+\beta^{2}=\alpha$ and once again the cauchy-Binet formula we finally get
$\det(K_{S})$ $=(1-\alpha)\sum_{\begin{subarray}{c}V\subset\{3,\cdots,l+2\}\\ |V|=|S|\end{subarray}}\det(P_{S,V})^{2}+\alpha\sum_{\begin{subarray}{c}V\subset[l+2]\\ |V|=|S|\end{subarray}}\det(P_{S,V})^{2}$
$=(1-\alpha)\det(K^{(2)}_{S})+\alpha\det(K^{(1)}_{S})$
$=(1-\alpha)\mathbb{P}(S\subset Y)+\alpha\mathbb{P}(S\subset X)=\mathbb{P}(S\subset Z),$
concluding the proof. ∎
Unfortunately, this result does not generalize well when $k\geqslant 2$. To illustrate the issue let $k\geqslant 2$ and $l=0$ so that the kernel $L$ is well-defined and of the form (3.6). We write $S_{i}=\begin{pmatrix}0&\mu_{i}\\ -\mu_{i}&0\end{pmatrix}$ to simplify the notations and we also write $p_{i}=\frac{\mu_{i}^{2}}{1+\mu_{i}^{2}}\in]0,1[$ and $q_{i}=1-p_{i}$. For odd-sized sets $S\subset[n]$ we have $\det(L_{S})=0$ since $L$ is skew-symmetric. When $S$ is of even size then the Pfaffian minor-summation formula [19] gives
$\det(L_{S})=\textnormal{Pf}(L_{S})^{2}=\left(\sum_{\begin{subarray}{c}A\subset[n]\\ |A|=|S|\end{subarray}}\textnormal{Pf}(D_{A})\det(P_{S,A})\right)^{2}$
For $\textnormal{Pf}(D_{A})$ to be non-zero then we obviously need to have $|A|\leqslant 2k$ and $A$ has to satisfy either $A\cap\{i,i+1\}=\emptyset$ or $\{i,i+1\}\subset A$ for any odd $i\leqslant 2k-1$. Now, for any set $A\subset[k]$ we write $\phi(A):=\cup_{i\in A}\{2i-1,2i\}$. Then,
$\frac{\det(L_{S})}{\det(I_{n}+L)}$ $=\prod_{i=1}^{k}(1+\mu_{i}^{2})^{-1}\left(\sum_{\begin{subarray}{c}A\subset[k]\\ |A|=|S|/2\end{subarray}}\prod_{i\in A}\mu_{i}\det(P_{S,\phi(A)})\right)^{2}$
$=\left(\sum_{\begin{subarray}{c}A\subset[k]\\ |A|=|S|/2\end{subarray}}\prod_{i\in A}\sqrt{p_{i}}\prod_{i\notin A}\sqrt{q_{i}}\det(P_{S,\phi(A)})\right)^{2}.$
When $k=1$ we recover the result of Proposition 3.7:
$\frac{\det(L_{S})}{\det(I_{n}+L)}=\left\{\begin{array}[]{l}p_{1}\det(P_{S,\{1,2\}})^{2}\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ |S|=2;\\ q_{1}\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ |S|=0;\end{array}\right.$
but when $k\geqslant 2$ then there is no clear expression for $\frac{\det(L_{S})}{\det(I_{n}+L)}$ as a mixture of probability distributions. The problem of adapting [18, Algorithm 4.4.2] to the case $k\geqslant 2$ thus remains an open problem.
4 Application to the simulation of attractive couplings of repulsive DPPs
We consider in this section the possible use of DPPs with nonsymmetric kernels to construct coupling of DPPs that model marked point processes with repulsion between points of the same mark and attraction between points of different marks. One such example is the ants dataset [16] in the spatstat R package [3] containing locations of nests of two species of ants. While ants of the same species tends to spread out their nests to avoid having to share resources, hence the repulsion in their nest location, it is suspected that one of the two species purposefully put their nest close to one of the other species since they eat dead insects, hence the attraction between the location of nests of different species.
4.1 DPP couplings
We recall a natural way of making couplings of determinantal point processes. To express such couplings we begin by pointing out that a coupling of two point processes on $[n]$ can be seen as a point process on $[2n]$ using the bijection
$\begin{array}[]{ccc}[2n]&\leftrightarrow&[n]\times[n]\\ X&\leftrightarrow&(X_{1},X_{2})\end{array}\leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \left\{\begin{array}[]{l}X_{1}=X\cap[n];\\ X_{2}=\big{\{}i-n,i\in X\cap\{n+1,\cdots,2n\}\big{\}}.\end{array}\right.$
As an abuse of notation, we thus write any DPP $X$ on $[2n]$ as a coupling $(X_{1},X_{2})$ of DPPs on $[n]$. Now, we consider two DPP kernels $K^{(1)},K^{(2)}\in\mathcal{M}_{n}(\mathbb{R})$ and a $2\times 2$ block matrix of the form
$\mathbb{K}=\begin{pmatrix}K^{(1)}&*\\ *&K^{(2)}\end{pmatrix}\in\mathcal{M}_{2n}(\mathbb{R})$ (4.1)
such that $\mathbb{K}$ is a DPP kernel. If $(X_{1},X_{2})\sim\textnormal{DPP}(\mathbb{K})$ then $X_{1}\sim\textnormal{DPP}(K^{(1)})$ and $X_{2}\sim\textnormal{DPP}(K^{(2)})$ by Proposition 1.4 (a). This way, we can write a coupling of two DPPs on $[n]$ as a DPP on $[2n]$. With a similar abuse of notation we also write $\mathbb{K}_{(S,T)}$ for the submatrix of $\mathbb{K}$ where we kept the rows and columns of $\mathbb{K}$ indexed by $S\cup\{i+n,i\in T\}$.
Such a setting has been considered before [14] but only using symmetric kernels. When $\mathbb{K}$ is symmetric then
$\mathbb{P}(i\in X_{1},j\in X_{2})-\mathbb{P}(i\in X_{1})\mathbb{P}(j\in X_{2})=-\mathbb{K}_{i,j+n}^{2}\leqslant 0$
and therefore this coupling only allows negative cross-correlations. This is why we suggest the possibility of using non-symmetric matrices in order to use this very natural way of constructing DPP couplings while allowing the coupling to have some positive cross-dependencies. While it is possible to find other ways of constructing couplings of DPPs, the fact that we can write the coupling as a DPP itself is important since it would allow us to use the nice statistical properties of DPPs such as estimation using their moments [24, 25], their likelihood [29, 11] and the various associated CLTs [4, 28, 9].
4.2 Examples of couplings
One of the earliest example of construction of a specific DPP coupling (using only symmetric DPP kernels) comes from [26] where the author considered a symmetric DPP kernel $K$ and the coupling with kernel
$\mathbb{K}=\begin{pmatrix}K&\sqrt{K(I_{n}-K)}\\ \sqrt{K(I_{n}-K)}&I_{n}-K\end{pmatrix}$
which has the nice property of having a deterministic number of points since $\mathbb{K}$ is a projection matrix. This coupling is also an important tool to show that DPPs with symmetric kernels are strongly Rayleigh [5].
In order to illustrate that we can use (4.1) to construct couplings with various degrees of repulsion and attraction we consider the following setting. Let $X$ be a DPP, we construct two new point processes $X_{1}$ and $X_{2}$ as subsets of $X$ the following way: for each point $i\in X$ let $B_{i}\sim b(\frac{p_{i}}{2-p_{i}})$, for some $p_{i}\in[0,1]$. If $B_{i}=1$ then $X$ gives the point $i$ to both $X_{1}$ and $X_{2}$, otherwise $X$ gives the point $i$ to either $X_{1}$ with probability $1/2$ or $X_{2}$ with probability $1/2$. As a result, we get that $X_{1},X_{2}\subset X$ and the higher the probabilities $p_{i}$ are the more points in common both $X_{1}$ and $X_{2}$ will get. We now show that $(X_{1},X_{2})$ can be written as a DPP coupling of the form (4.1).
Proposition 4.1.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ be a DPP kernel and let $p=(p_{1},\cdots p_{n})\in[0,1]^{n}$ such that $D(2\cdot 1_{n}-p)K$ is a DPP kernel. Let $X\sim\textnormal{DPP}(D(2\cdot 1_{n}-p)K)$, let $B_{1},\cdots,B_{n}$ be Bernoulli random variables with $B_{i}\sim b(\frac{p_{i}}{2-p_{i}})$ and $B^{\prime}_{1},\cdots,B^{\prime}_{n}$ be Bernoulli random variables with $B^{\prime}_{i}\sim b(1/2)$. All $B_{i}$, $B^{\prime}_{i}$ and $X$ are assumed to be mutually independent. We construct $X_{1}$ and $X_{2}$ the following way:
$\begin{array}[]{l}X_{1}=\{i\in X,\leavevmode\nobreak\ B_{i}=1\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ B_{i}=0\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B^{\prime}_{i}=1\},\\ X_{2}=\{i\in X,\leavevmode\nobreak\ B_{i}=1\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ B_{i}=0\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B^{\prime}_{i}=0\}.\end{array}$
Then, $(X_{1},X_{2})\sim\textnormal{DPP}(\mathbb{K})$ where
$\mathbb{K}=\begin{pmatrix}K&K-D(p)\\ K&K\end{pmatrix}.$
Proof.
Let $S,T\subset[n]$. We write $A=S\cap T$ and decompose $S$ and $T$ into $S=S^{\prime}\cup A$ and $T=T^{\prime}\cup A$ with $A\cap S^{\prime}=A\cap T^{\prime}=\emptyset$. By definition of $X_{1}$ and $X_{2}$ we have
$\mathbb{P}(S\subset X_{1},T\subset X_{2})$
[…]
×$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \times\mathbb{P}(\forall i\in T^{\prime},\leavevmode\nobreak\ B_{i}=1\leavevmode\nobreak\ \mbox{or}\leavevmode\nobreak\ B_{i}=0\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ B^{\prime}_{i}=0)\mathbb{P}(\forall i\in A,\leavevmode\nobreak\ B_{i}=1)$
[…]
Moreover, with the right permutation of rows and columns we can write
$\det(\mathbb{K}_{(S,T)})$ $=\det\begin{pmatrix}K_{S^{\prime}}&K_{S^{\prime},A}&K_{S^{\prime},A}&K_{S^{\prime},T}\\ K_{A,S^{\prime}}&K_{A}&K_{A}-D(p)_{A}&K_{A,T}\\ K_{A,S^{\prime}}&K_{A}&K_{A}&K_{A,T}\\ K_{T^{\prime},S^{\prime}}&K_{T^{\prime},A}&K_{T^{\prime},A}&K_{T^{\prime}}\\ \end{pmatrix}$
$=\det\begin{pmatrix}K_{S^{\prime}}&K_{S^{\prime},A}&K_{S^{\prime},A}&K_{S^{\prime},T}\\ K_{A,S^{\prime}}&K_{A}&K_{A}-D(p)_{A}&K_{A,T}\\ 0&0&D(p)_{A}&0\\ K_{T^{\prime},S^{\prime}}&K_{T^{\prime},A}&K_{T^{\prime},A}&K_{T^{\prime}}\\ \end{pmatrix}$
$=\det(D(p)_{A})\det(K_{S\cup T}).\qed$
Note that when $p=1_{n}$ in Proposition 4.1 then $X_{1}$ and $X_{2}$ are constructed by splitting each point of $X$ into either $X_{1}$ or $X_{2}$ with probability $\frac{1}{2}$. It corresponds to a very repulsive coupling since $X_{1}\cap X_{2}=\emptyset$ almost surely. This particular case matches the setting of [1] and we recover the following result that was pointed out in [14]:
Corollary 4.2.
Let $K\in\mathcal{M}_{n}(\mathbb{R})$ such that $2K$ is a DPP kernel. Let $X\sim\textnormal{DPP}(2K)$. We randomly split $X$ into two disjoint subsets $X_{1}$ and $X_{2}$ such that for each $i\in X$ we choose independently and with probability $1/2$ whether $i\in X_{1}$ or $i\in X_{2}$. Then, $(X_{1},X_{2})\sim\textnormal{DPP}(\mathbb{K})$ where
[…]
When $p=0_{n}$ in Proposition 4.1 then $X_{1}=X_{2}=X$, corresponding to the most attractive coupling possible of two DPPs, and we get the following result:
Corollary 4.3.
Let $X\sim\textnormal{DPP}(K)$ where $K\in\mathcal{M}_{n}(\mathbb{R})$ is a DPP kernel. Then $(X,X)\sim\textnormal{DPP}(\mathbb{K})$ where
$\mathbb{K}=\begin{pmatrix}K&K-I_{n}\\ K&K\end{pmatrix}.$
Both cases shows that the setting of (4.1) allows to construct both very repulsive and very attractive couplings of two DPPs with the same distribution. Note that (4.1) also allows the construction of an independent couplings of any DPP with kernel $K^{(1)}$ and any DPP with kernel $K^{(2)}$ since such a coupling can be written as a DPP with kernel $\mathbb{K}=\begin{pmatrix}K^{(1)}&*\\ 0&K^{(2)}\end{pmatrix}$.
4.3 Construction of an attractive coupling of repulsive DPPs
In order to construct an attractive coupling of two repulsive DPPs we consider $K$ to be a symmetric matrix with eigenvalues in $[0,1]$ (and thus a DPP kernel) and we construct a coupling $(X_{1},X_{2})\sim\textnormal{DPP}(\mathbb{K})$ on $[2n]$ where $\mathbb{K}$ is of the form
$\mathbb{K}=\begin{pmatrix}K&A\\ B&K\end{pmatrix}.$ (4.2)
We then have $X_{1}\sim\textnormal{DPP}(K)$, $X_{2}\sim\textnormal{DPP}(K)$ and since $K$ is symmetric then $X_{1}$ and $X_{2}$ are both repulsive point processes. Now, the cross dependency between $X_{1}$ and $X_{2}$ can be observed using the quantities
$\mathbb{P}(i\in X_{1},j\in X_{2})-\mathbb{P}(i\in X_{1})\mathbb{P}(j\in X_{2})=-A_{i,j}B_{j,i}.$
In order to have positive cross correlations we thus need to assume at minimum that $A_{i,i}B_{i,i}\leqslant 0$ for all $i\in[n]$. A natural way of satisfying this assumption is to take $A=-B^{T}$ although Proposition 1.5 gives
$(X_{1},X_{2})\sim\textnormal{DPP}\left(\begin{pmatrix}K&A\\ -A^{T}&K\end{pmatrix}\right)\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ (X_{1},X_{2}^{c})\sim\textnormal{DPP}\left(\begin{pmatrix}K&A\\ A^{T}&I_{n}-K\end{pmatrix}\right).$
Thus, the distribution of $(X_{1},X_{2}^{c})$ is a DPP with a symmetric kernel and we recover a well known setting [14]. Unfortunately, this setting is too restrictive. It does not cover the couplings presented in Proposition 4.3 and when using it in our simulations we were only able to use such kernels to generate DPP couplings with barely any attraction. Another natural way of constructing positive cross correlations is to take $A$ (resp. $B$) to be a symmetric positive (resp. negative) semidefinite matrix. By additionally assuming that $A,B$ and $K$ commute, and are therefore simultaneously diagonalizable, and then using Proposition 2.4 we can choose $A$ and $B$ such that $\mathbb{K}$ is a DPP kernel with the following result.
Proposition 4.4.
Let $K\in\mathcal{S}^{+}_{n}(\mathbb{R})$ with eigenvalues $\lambda=(\lambda_{1},\cdots,\lambda_{n})\in[0,1]^{n}$ and write $K=PD(\lambda)P^{T}$ the spectral decomposition of $K$. Let $\mu=(\mu_{1},\cdots,\mu_{n})\in\mathbb{R}_{+}^{n}$ and $\nu=(\nu_{1},\cdots,\nu_{n})\in\mathbb{R}_{+}^{n}$ such that
$(\lambda_{i}-1/2)^{2}+\frac{1}{2}\left(\mu_{i}^{2}+\nu_{i}^{2}+|\mu_{i}-\nu_{i}|\sqrt{4\left(\lambda_{i}-\frac{1}{2}\right)^{2}+(\mu_{i}+\nu_{i})^{2}}\right)\leqslant\frac{1}{4}$ (4.3)
and define $A=PD(\mu)P^{T}\in\mathcal{S}^{+}_{n}(\mathbb{R})$ and $B=-PD(\nu)P^{T}\in\mathcal{S}^{+}_{n}(\mathbb{R})$. Then $\mathbb{K}=\begin{pmatrix}K&A\\ B&K\end{pmatrix}$ is a DPP kernel.
Proof.
We can write $\mathbb{K}$ as
$\mathbb{K}$ $=\begin{pmatrix}P&0\\ 0&P\end{pmatrix}\begin{pmatrix}D(\lambda)&D(\mu)\\ -D(\nu)&D(\lambda)\end{pmatrix}\begin{pmatrix}P^{T}&0\\ 0&P^{T}\end{pmatrix}$
$=\frac{1}{2}\left(I_{2n}-\mathbb{M}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{M}=-2\begin{pmatrix}P&0\\ 0&P\end{pmatrix}\begin{pmatrix}D(\lambda-\frac{1}{2}1_{n})&D(\mu)\\ -D(\nu)&D(\lambda-\frac{1}{2}1_{n})\end{pmatrix}\begin{pmatrix}P^{T}&0\\ 0&P^{T}\end{pmatrix}.$
We then have
$\mathbb{M}\mathbb{M}^{T}=4\begin{pmatrix}P&0\\ 0&P\end{pmatrix}\begin{pmatrix}D((\lambda-\frac{1}{2}1_{n})^{2}+\mu^{2})&D((\mu-\nu)(\lambda-\frac{1}{2}1_{n}))\\ D((\mu-\nu)(\lambda-\frac{1}{2}1_{n}))&D((\lambda-\frac{1}{2}1_{n})^{2}+\nu^{2})\end{pmatrix}\begin{pmatrix}P^{T}&0\\ 0&P^{T}\end{pmatrix}$
we can then deduce that the $2n$ eigenvalues of $\mathbb{M}\mathbb{M}^{T}$ (and thus the singular values squared of $\mathbb{M}$) are equal to four times the two eigenvalues of the $n$ matrices
$\begin{pmatrix}(\lambda_{i}-1/2)^{2}+\mu_{i}^{2}&(\mu_{i}-\nu_{i})(\lambda_{i}-1/2)\\ (\mu_{i}-\nu_{i})(\lambda_{i}-1/2)&(\lambda_{i}-1/2)^{2}+\nu_{i}^{2}\end{pmatrix}$
for each $i\in[n]$. It is straightforward to show that these eigenvalues are
$(\lambda_{i}-1/2)^{2}+\frac{1}{2}\left(\mu_{i}^{2}+\nu_{i}^{2}\pm|\mu_{i}-\nu_{i}|\sqrt{4\left(\lambda_{i}-\frac{1}{2}\right)^{2}+(\mu_{i}+\nu_{i})^{2}}\right).$
If these quantities are all $\leqslant 1/4$ then Proposition 2.4 shows that $\mathbb{K}$ is a DPP kernel. ∎
Note that since the first term of (4.3) has values in $[0,\frac{1}{4}]$ and the second term can be made arbitrarily small by choosing $\mu_{i}$ and $\nu_{i}$ close to 00 then it is always possible to find some $\mu$ and $\nu$ satisfying (4.3).
4.4 Simulation of non-symmetric DPPs
As mentioned in Section 3.4, the main method used to simulate both discrete and continuous DPPs with symmetric kernels uses a symmetric version of Proposition 3.4 to turn the DPP kernel into a projection matrix then simulate projection DPPs with [18, Algorithm 4.4.2]. While this is the only known method used to perfectly simulate continuous DPPs, other algorithms have been developed to simulate discrete DPPs, even with non-symmetric kernels. Since we were not able to modify [18, Algorithm 4.4.2] for DPPs with non-symmetric kernel we thus relied on using the generic algorithm of [30] for our simulations. We also mention [15, 23] for different algorithms that can be used to simulate discrete DPPs with non-symmetric kernels.
4.5 Numerical results
[…]
(a) Regular grid of $31\times 31$ points
(b) Subset of the random grid chosen by a DPP
Figure 2: Example of a simulation of a DPP on a $31\times 31$ regular grid of $[0,1]^{2}$ with a symmetric kernel $K$ satisfying $K_{i,j}=0.02\exp\left(-\frac{\|P_{i}-P_{j}\|_{2}^{2}}{0.018}\right).$
We then show in Figures 3 and 4 two attractive couplings of two repulsive DPPs with the same kernel on a regular grid of $[0,1]^{2}$ with $31\times 31$ points. In both cases we simulate a coupling $(X_{1},X_{2})$ with kernel of the form (4.2). For Figure 3 we chose $K_{i,j}=0.02\exp\left(-\frac{\|P_{i}-P_{j}\|_{2}^{2}}{0.018}\right)$, called a Gaussian kernel, and for Figure 4 we chose $K_{i,j}=0.02\left(1+\left(\frac{\|y-x\|}{0.075}\right)^{2}\right)^{-1.1}$, called a Cauchy kernel. In both cases, the matrices $A$ and $B$ were chosen using the method described in Proposition 4.4 with all $\mu_{i}$ and $\nu_{i}$ chosen randomly among the eigenvalues satisfying (4.3). Figures 3a and 4a both shows a simulation of the coupling $(X_{1},X_{2})$. In order to illustrate the attraction between $X_{1}$ and $X_{2}$ we begin from the simulation we obtained of $(X_{1},X_{2})$ then we "forget" the value of $X_{2}$ and we compute the probability that each point $P_{i}$ is in $X_{2}$ given the value of $X_{1}$ we obtained. We then present these values as a heatmap in Figures 3b and 4b.
(a) Simulation of the coupling of $X_{1}$ (squares) and $X_{2}$ (circles).
(b) Inclusion probability for $X_{2}$ given the simulation of $X_{1}$ obtained in Figure 3a
Figure 3: Illustration of an attractive coupling $(X_{1},X_{2})$ of two DPPs with a Gaussian Kernel.
(a) Simulation of the coupling of $X_{1}$ (squares) and $X_{2}$ (circles). Points in $X_{1}\cap X_{2}$ are displayed as triangles.
(b) Inclusion probability for $X_{2}$ given the simulation of $X_{1}$ obtained in Figure 4a
Figure 4: Illustration of an attractive coupling $(X_{1},X_{2})$ of two DPPs with a Cauchy Kernel.
Both simulations in Figure 3 and 4 display the behaviour we intended with points of $(X_{1},X_{2})$ being more likely to be close to each other than if $X_{1}$ and $X_{2}$ were simulated independently. Both examples also shows different kind of attraction with Figure 3 showing a small attraction with a long range, leading to points in $X_{1}$ and $X_{2}$ falling roughly in similar places, whereas Figure 3 show a strong attraction but with a very short range, leading to most points in $X_{1}$ and $X_{2}$ behaving like independent DPPs except for a few points that falls very close to each other or at the exact same location. These examples thus illustrates the possibility to use DPPs with non-symmetric kernels in order to construct coupling with varying degrees of attraction.
Of course, these examples are mostly a proof of concept. A lot more work is needed to see if there is a proper way to, at minimum, control the strength and range of attraction in the coupling as well as showing if this model is viable in statistical applications. But these tasks are outside the scope of this paper and they remained to be explored in future works.
5 Discussion and open problems
We showed in this paper the close link between the theory of $P_{0}$ matrices and the theory of DPPs, both area of Mathematics having some results being almost equivalent despite being expressed in a seemingly different setting. We used these close links to adapt some of the properties of $P_{0}$ matrices to describe the properties of DPPs with non-symmetric kernels. We focused in particular on DPPs with kernels of the form $\frac{1}{2}(I_{n}-M)$ with $\|M\|_{2}\leqslant 1$ that generalize symmetric kernels, are simple to construct and for whom we adapted some of the common properties of DPPs with symmetric kernels. We finally illustrated a possible use of DPPs with generic kernels for building models of marked point processes with repulsion between points of the same mark and attraction between points of different marks.
There is obviously still a lot more work to be done in order to deepen our understanding of the theory of DPPs so we finish by mentioning a non exhaustive list of important properties of DPPs with symmetric kernels whose potential generalization to non-symmetric kernels would be interesting to be further investigated.
Continuous setting
A lot of applications of DPPs, including the original purpose of modeling fermions [27], consider a continuous setting, meaning that DPPs are locally finite random subsets of a general Hilbert set $H$, usually $\mathbb{R}^{d}$. The symmetric kernel $K$ of finite DPPs is then replaced by a locally square integrable, locally of trace class hermitian operator $\mathcal{K}:H^{2}\rightarrow\mathbb{R}$ with eigenvalues in $[0,1]$. The immediate questions arising from the results in Sections 2 and Sections 3 is if it is possible to replace the assumption that $\mathcal{K}$ is hermitian with a continuous analogous to Proposition 2.4.
Negative association
All DPPs with symmetric kernels satisfy a strong dependency property called negative association [5, 26] that extends to continuous DPPs with numerous applications [12, 28]. Since negative association implies pairwise repulsion then a necessary condition for a DPP with kernel $K$ to be negatively associated is that $K_{i,j}K_{j,i}\geqslant 0$ for any $i,j\in[n]$ which is obviously not satisfied by all generic DPP kernels. The question naturally arising from this conclusion is what condition on $K$ is needed for the associated DPP to be negatively associated? In fact, does a negatively associated DPP that does not have the same distribution as a DPP with symmetric kernel even exist?
Stochastic dominance
Consider two symmetric DPP kernels $K$ and $K^{\prime}$. If $K^{\prime}-K$ is a positive semidefinite matrix then the DPP with kernel $K^{\prime}$ stochastically dominates the DPP with kernel $K$ [26, 5]. A similar result also exists for stationary continuous DPPs [13]. This is an important property with various applications for continuous DPPs such as describing their reach of repulsion [2] or the size of their Voronoï cells [13]. Now, for any general DPP kernel $K$ and $p\in[0,1]$ the DPP with kernel $K$ stochastically dominates the DPP with kernel $pK$ by Proposition 1.4 (b) even is $K-pK$ is not a positive semidefinite matrix. One can therefore wonder if there is a simple condition on a pair of kernel $(K,K^{\prime})$ that generalizes the symmetric case and indicates whether the DPP with kernel $K$ is stochastically dominated by the DPP with kernel $K^{\prime}$?